### v1

# The BPgWSP test: a Bayesian Weibull Shape Parameter signal detection test for adverse drug reactions

Abstract
We develop a Bayesian Power generalized Weibull shape parameter (PgWSP) test as statistical method for signal detection of possible drug-adverse event associations using electronic health records for pharmacovigilance. The Bayesian approach allows the incorporation of prior knowledge about the likely time of occurrence along time-to-event data. The test is based on the shape parameters of the Power generalized Weibull (PgW) distribution. When both shape parameters are equal to one, the PgW distribution reduces to an exponential distribution, i.e. a constant hazard function. This is interpreted as no temporal association between drug and adverse event. The Bayesian PgWSP test involves comparing a region of practical equivalence (ROPE) around one reflecting the null hypothesis with estimated credibility intervals reflecting the posterior means of the shape parameters. The decision to raise a signal is based on the outcomes of the ROPE test and the selected combination rule for these outcomes.
The development of the test requires a simulation study for tuning of the ROPE and credibility intervals to optimize specificity and sensitivity of the test. Samples are generated under various conditions, including differences in sample size, prevalence of adverse drug reactions (ADRs), and the proportion of adverse events. We explore prior assumptions reflecting the belief in the presence or absence of ADRs at different points in the observation period. Various types of ROPE, credibility intervals, and combination rules are assessed and optimal tuning parameters are identified based on the area under the curve. The tuned Bayesian PgWSP test is illustrated in a case study in which the time-dependent correlation between the intake of bisphosphonates and four adverse events is investigated.
Key words:
signal detection, time-to-event models, Bayesian test, generalized Weibull distribution, pharmacology, adverse drug reactions
1 Introduction
Making assumptions about the generating process of an outcome in the development of a test of association can improve the performance of this test in terms of power, specificity, and sensitivity by enabling the use of more statistical information included in the data. In the field of pharmacovigilance (surveillance of occurrence of adverse events (AEs) related to prescribed drugs (Meyboom et al.,, 1997), this approach has been used by Cornelius and colleagues (Cornelius et al.,, 2012; Sauzet and Cornelius,, 2022) to develop a family of signal detection tests (test of association between the occurrence of an AE and a prescribed drug) which is based on an assumption about time dependence on drug prescription. More specific, Cornelius et al., (2012) proposed tests aiming for detection of a non-constant hazard function by modeling the time to occurrence of event from the time of prescription to flag a possible drug-event association. However, this is not the only type of information about a generating process that can be incorporated into a test. Prior knowledge regarding the distribution parameters of the generating process (e.g. mean time of occurrence of an event) can be incorporated in the development of a test using a Bayesian approach. The aim of this article is to combine both approaches by developing a Bayesian test to detect non-constant hazard functions for a time-dependent process incorporating prior knowledge on the distribution of the time to the occurrence of events. While the principle is rather general, we present this test in the context of pharmacovigilance which we briefly describe below, and review some existing methods of signal detection that assume a temporal relationship or use a Bayesian approach.
After release of a drug on the market, pharmacovigilance monitors the occurrences and changes of known adverse drug reactions (ADRs) as well as aims to detect new ADRs in the population. An ADR is an unintended, harmful response to a medicine (EMA,, 2024).
The investigation of a harm profile of a drug consists of multiple steps: screening the range of AEs that occur along prescribed drugs, performing a signal detection test to investigate the association for a drug-AE pair and, in case a signal is raised, deeper investigation of any causal relationship between the two which involves a complex epidemiological analysis (Meyboom et al.,, 1997).
Longitudinal electronic health records (EHRs) gathered by health insurance companies and health regulatory organizations such as the Sentinel system established by the FDA (Food and Drug Administration,, 2022) provide regularly collected longitudinal data about diagnoses, treatments, and medication prescribed by physicians, as well as AEs (Suling and Pigeot,, 2012; Sturkenboom and Schink,, 2021).
With the collection of EHR data, statistical methods to exploit them have emerged. These include among others extensions of disproportionality approaches comparing rates of occurrence between a single drug and the rest of the dataset (Zorych et al.,, 2013) for application to longitudinal data (Schuemie,, 2011), temporal pattern analysis (Norén et al.,, 2010), exposure models (Dijkstra et al.,, 2024) and more (see for example the review by Coste et al., (2023)).
As mentioned above Cornelius et al., (2012) and Sauzet and Cornelius, (2022) considered the temporal dynamics of the risk to experience a specific AE in their family of Weibull shape parameter (WSP) family of tests. For this a Weibull type distribution parametric hazard function is fitted to the data. A signal is raised following the rejection of a null-hypothesis assuming a constant hazard which means that a sufficiently large change in risk over time can be interpreted as a signal. Unlike disproportionality methods, no group comparisons are performed.
Bayesian statistical methods for signal detection have also been developed. But methods like the Empirical Bayes Gamma Mixture/ Gamma-Poisson shrinker or the Bayesian confidence propagation neural network are Bayesian extensions of disproportionality measures used to obtain signals from spontaneous reporting databases (Bansal et al.,, 2024) which do not consider the time of the event occurrence.
The objective of this work is to combine an assumption of temporal process with the incorporation of prior knowledge about distribution parameters of this process in the development of Bayesian Weibull shape parameter tests. To do so, we will generalize the proposed power generalized Weibull shape parameter (PgWSP) test by Sauzet and Cornelius, (2022) into a Bayesian framework to develop and tune an HDI+ROPE type test of signal detection.
The remaining article is structured as follows. In section 2 the theoretical concept of the Bayesian PgWSP (BPgWSP) test is presented. In section 3, we conduct a simulation study to tune the parameters defining the BPgWSP test in a chosen framework of simulation scenarios. Finally, in section 4, we apply our tuned BPgWSP test to data on women prescribed with bisphosphonates from the The Health Improvement Network (THIN) (THIN,, 2024) database in a case study.
2 Methodological concept for the Bayesian signal detection test
Figure 1: Methodological concept for the signal detection test.
The principle of non-constant hazard signal detection tests is built on the assumption that the risk of an AE is time-dependent relative to the prescription of the drug under investigation.
The parametric distribution is defined by a parameter vector $\Theta$. We make use of a subset of parameters $\mathcal{N}\subset\Theta$ on which to define the null-hypothesis of constant hazard and develop the Bayesian signal detection test based on an updated distribution for $\mathcal{N}$ using the HDI+ROPE approach (see sec. 2.2). For this we will perform the three steps below:
  1. 1.
    present the power generalized Weibull (PgW) distribution which we will use in this paper and formulate the null-hypothesis,
  2. 2.
    describe the BPgWSP test using the HDI+ROPE approach,
  3. 3.
    set the prior specification and HDI+ROPE tuning parameters for the BPgWSP test by performing a simulation study (see sec. 3).
2.1 Power generalized Weibull distribution:
The survival function of the PgW distribution is given by:
$S(t)=\text{exp}\left\{1-\left[1+\left(\frac{t}{\theta}\right)^{\nu}\right]^{\frac{1}{\gamma}}\right\}.$ (1)
The hazard function is given by:
$h(t)=\frac{\nu}{\gamma\theta^{\nu}}\cdot t^{\nu-1}\cdot\left[1+\left(\frac{t}{\theta}\right)^{\nu}\right]^{\frac{1}{\gamma}-1},$ (2)
(see Bagdonavicius and Nikulin, (2001, p.8); Nikulin et al., (2016, p.8)). The PgW distribution has three parameters $\Theta_{PgW}=\{\theta,\nu,\gamma\}$ two of which are shape parameters such that $\mathcal{N}_{PgW}=\{\nu,\gamma\}$.
Figure 2: Hazard functions obtained from power generalized Weibull distributions with a) constant ($\theta=2,\nu=1,\gamma=1$) , b) increasing ($\theta=8,\nu=3,\gamma=1$), c) decreasing ($\theta=1.5,\nu=0.6,\gamma=1$), d) unimodal ($\theta=0.5,\nu=5,\gamma=10$), e) bathtub ($\theta=30,\nu=0.9,\gamma=0.2$) form.
The two shape parameters allow for a range of hazard function forms including constant, monotoneously in- or decreasing and unimodal or bathtub shaped (see Figure 2).
A constant hazard is given if and only if both shape parameters are equal to one so that the null-hypothesis of constant hazard can be formulated as:
$H_{0}:\nu=1\text{ and }\gamma=1.$ (3)
2.2 Development of Bayesian signal detection test using the HDI+ROPE approach
Prior and posterior distribution
The prior distribution of $\Theta$ should reflect whether the AE of interest is suspected to be an ADR (non-constant hazard) and if so, when it is expected to occur (specifying $\mathbf{E}[t]$) a priori by medical experts. This can be translated into the corresponding expected hazard form (comp. Figure 2) parameterized in terms of prior mean values for the parameters in $\Theta$. The prior distribution of the scale parameter reflects the AE background rate. The priors of the shape parameters $\mathcal{N}\subset\Theta$ determine the constancy or non-constancy and general trend of the hazard over time.
Based on the specified prior means (and a chosen standard deviation) as well as a distributional choice (eg. a Lognormal distribution for each parameter in $\Theta$), the prior distribution $p_{0}(\Theta)$ will be updated with the data available. For time-to-event data the posterior distribution $p_{1}(\Theta|t)$ is proportional to
$p_{1}(\Theta|t)\propto p_{0}(\Theta)\cdot\mathcal{L}(t|\Theta)$ (4)
where
$\mathcal{L}(t|\Theta)=\prod_{i=1}^{N}S(t_{i})^{1-d_{i}}\cdot f(t_{i})^{d_{i}}$ (5)
with $\mathcal{L}(t|\Theta)$ being the likelihood function in the time-to-event framework such that all subjects $i$ not having an AE during the observation time (denoted by $d_{i}=0$) are considered in terms of the survival function $S(t_{i})$ and all subjects $i$ experiencing an AE (denoted by $d_{i}=1)$ are taken into account in terms of the density function $f(t_{i})=S(t_{i})\cdot h(t_{i})$.
HDI+ROPE method
Based on the posterior distribution of $\mathcal{N}$, the HDI+ROPE test (Kruschke,, 2015, 2018) applied to each shape parameter and combination of all single shape parameter test results provide two possible outcomes: either we lean towards the null-hypothesis of constant hazard or the alternative hypothesis of a non-constant hazard. The test setup depends on
  • the region of practical equivalence (ROPE) to the null value, in our case one, typically defined as an interval around the null value and
  • a credibility interval (CI) based on posterior distribution parameters in $\mathcal{N}$, here either a highest density interval (HDI) as proposed by Kruschke, (2015) or an equal tailed interval (ETI). See sec. 3.3 for the formal definition of the ETI and HDI.
Figure 3: HDI+ROPE decision rules based on graphic from Kruschke, (2018, Fig. 1).
In our setting, the HDI+ROPE test is applied to each parameter $\nu\in\mathcal{N}$ and provides one of the following interim results (compare to Figure 3):
  • If the CI lies completely within the ROPE, then the single parameter null-hypothesis
    $H_{0}(\nu):\nu=1$
    is accepted,
  • if ROPE and CI do not intersect at all, then the single parameter null-hypothesis is rejected,
  • if ROPE and CI partly overlap, then no decision is made (undecided).
Given the single parameter test outputs, the combined test result is concluded according to a combination rule, for instance the condition directly implied by Equation 3:
  • If the HDI+ROPE accepts the null-hypothesis for all parameters in $\mathcal{N}$, then accept the null-hypothesis of constant hazard.
  • Else, reject the null-hypothesis of constant hazard and raise a signal.
To complete the test development, we need to make decisions about the distributional choice for the prior, the position and width of the ROPE and the CI of the BPgWSP test and the combination rule. This will be done in the following section.
3 Simulation study for test development and tuning
A method to finalize the BPgWSP test theoretically described in section 2 is to perform a simulation study. By applying a range of prior distributions, choices for position and width of the ROPE, position and width of the CI and the combination rules, we make decisions about the optimal tuning parameters for the BPgWSP test such that a performance parameter based on sensitivity and specificity of the test is maximal across a range of sample scenarios.
3.1 Data generating process
The set of simulation scenarios is determined by the
  • censoring time (365 days) which sets the end of the observation period,
  • sample size $N$ (500, 3000, 5000),
  • backround rate (0.1) which specifies the prevalence in the population,
  • ADR rate which is the proportion of the background rate (0, 0.5, 1),
  • expected event-time of the ADR if the ADR rate is above zero (1st, 2nd, 3rd quarter of observation period),
  • relative standard deviation of the ADR event-time (0.05).
The background observations are sampled using a Binomial distribution with a probability equal to the specified background rate
$n^{br}\sim Bin(n=N,p=br).$
Corresponding event-times are sampled using a Uniform distribution over the study period ranging from day one to day $365$.
$t^{br}$ $\sim Unif(1,365)$ (6)
The number of events induced by the drug intake is sampled from a Binomial distribution with a probability equal to the background rate times the ADR rate.
$n^{ADR}\sim Bin(n=N,p=br\cdot adr)$ (7)
The event-times for the ADR cases are generated from a Normal distribution with an expected event-time $\mathrm{E}[t]$ at the first quarter, half, and third quarter of the study period and a standard deviation of $0.05$ times the censoring time at 365 days.
$t^{ADR}$ $\sim N(\mathrm{E}[t],0.05\cdot 365)$ (8)
The sample is filled up to size $N$ with censored observations. They have the AE status zero and time point $365$ equaling the end of the study period.
3.2 Prior specifications for the Bayesian model
Figure 4: Hazard functions obtained under the prior belief that the expected event-time is at the first, second, or third quarter of the observation period (left) or that the AE of interest is no ADR (right).
prior belief about prior means
expected event-time $(\bar{\theta}_{0},\bar{\nu}_{0},\bar{\gamma}_{0})$
none $(1,1,1)$
1st quarter of OP (91 days) $(1,0.207,1)$
2nd quarter of OP (183 days) $(20,5.5,14)$
3rd quarter of OP (274 days) $(300,4,1)$
Table 1: Mean of prior distribution under implemented prior beliefs about expected event-time.
Four different prior beliefs are implemented. An ADR is expected to occur with higher risk either around the first, second or third quarter of the observation period or not expected at all. These assumptions are represented by hazard functions of the distribution shown in Figure 4 based on chosen prior mean values for the scale $\theta$, and shapes $\nu$ and $\gamma$ listed in Table 1. The standard deviation of the parameters’ prior distributions is set to 10.
scale $\theta$ shape $\nu$ powershape $\gamma$ abbreviation
fixed to prior mean Gamma Gamma fix-gam-gam
Gamma Gamma Gamma gam-gam-gam
fixed to prior mean Lognormal Lognormal fix-log-log
Lognormal Lognormal Lognormal log-log-log
Table 2: Prior distributional assumptions for the scale $\theta$, and shapes $\nu$ and $\gamma$ of the PgW distribution.
For practical reasons, scale and shape parameters are assumed to be independent. By definition all parameters of the PgW distribution are positive. Four prior distributional choices matching these conditions are listed in Table 2. All parameters’ univariate priors follow either the Gamma or Lognormal distribution. In two of the four prior distribution choices, the scale parameter of the posterior remains equal to the prior (i.e. will not be updated).
For instance, the PgW model with log-log-log prior is be specified as follows:
tsimilar-$t\sim$ $PgW(\theta,\nu,\gamma)$ (9)
$\theta_{0}\sim Lognormal(\mu_{0}^{\theta},\sigma_{0}^{\theta})$ (10)
$\nu_{0}\sim Lognormal(\mu_{0}^{\nu},\sigma_{0}^{\nu})$ (11)
$\gamma_{0}\sim Lognormal(\mu_{0}^{\gamma},\sigma_{0}^{\gamma})$ (12)
where the prior meta parameters of the Lognormal distribution are the logarithm of the location $\mu$ and the logarithm of the scale $\sigma$ which can be calculated from the prior mean (specified in Table 1 and standard deviation (10) as
$\mu_{0}^{\theta}$ $=\log(\bar{\theta}_{0})-\log\left(\frac{sd(\theta_{0})^{2}}{\bar{\theta}_{0}^{2}}+1\right)$ (13)
$\sigma_{0}^{\theta}$ $=\sqrt{\log(\frac{sd(\theta_{0})^{2}}{\bar{\theta}_{0}^{2}}+1)}$ (14)
for $\theta$ and analogously for $\nu$ and $\gamma$.
A posterior estimate is obtained by Markov chain Monte Carlo (MCMC) sampling. We draw a posterior sample of size 40000 obtained from four chains of length 10000 after a burn-in phase of 1000 draws each.
Each combination of sample scenario and prior specification leads to a total of 336 different simulation settings – 288 with a positive ADR rate and 48 controls to match in with only background events for the simulation study. Each simulation scenario’s data generation and posterior estimation is repeated 100 times.
3.3 Specification of the HDI+ROPE test
The HDI-ROPE test approach is specified with respect to the
  • ROPE position and width,
  • posterior CI position and width,
  • combination of the single test outputs of both shape parameters.
We specify the ROPE for each shape parameter as equal-tailed interval (ETI)
$[q_{(1-\alpha)/2},q_{(1+\alpha)/2}]$ (15)
based on the quantiles of the shape parameters’ prior distributions under the null-hypothesis at credibility level $\alpha$.
We specify the posterior CI for each shape parameter either as an ETI calculated from the empirical quantiles of the shape parameters’ posterior samples at credibility level $\alpha$ or as an HDI at credibility level $\alpha$ calculated from the posterior sample. An HDI at level $\alpha$ is an interval
[…]
that is with boundaries such that all values inside the HDI are have higher density than values outside the HDI and simultaneously encompasses $(1-\alpha)\cdot 100\%$ of the probability mass (Kruschke,, 2015, p.87).
To control the width of the ROPE as well as the posterior CIs, we apply credibility levels $\alpha$ ranging from $0.5,0.55,...0.95$.
Three different combination rules of the single shape parameters’ test outcomes are defined in Table 3. Option 1 implements the hypothesis formulated in Equation 3. A signal is raised if at least one shape parameter significantly differs from one indicating a non-constant hazard function. Option 2 raises a signal in fewer cases, namely if at least one shape parameter’s ROPE test is positive and the other undecided. Option 3 raises a signal if and only if both single shape parameter tests are positive.
The simulation study is implemented in the statistical software environment R (Team,, 2022) including the packages rstan (Stan Development Team,, 2024) for the Bayesian model setup and posterior sampling, HDInterval (Meredith and Kruschke,, 2022) for calculation of posterior CIs, and BWSPsignal (available on https://github.com/julia-dyck/BWSPsignal) containing functions specifically written for the BPgWSP test development and simulation setup.
interim result combination rule
option 1 option 2 option 3
for $\nu$ for $\gamma$ (acc. to Eq. 3) (reserved) (very reserved)
rejected rejected signal signal signal
accepted rejected signal - -
rejected accepted signal - -
accepted accepted - - -
undecided rejected signal signal -
undecided accepted - - -
rejected undecided signal signal -
accepted undecided - - -
undecided undecided signal - -
Table 3: Three options for the combination of the single test outcomes for shape parameter $\nu$ and powershape parameter $\gamma$. Option 1 raises a signal if at least one parameter significantly differs from one. Option 2 raises a signal in fewer cases. Option 3 raises a signal only if both single parameter tests are positive.
3.4 Evaluation criteria for the simulation results
The evaluation of the simulation study is performed in three steps. The first step consists of selecting a combination of prior distributions (see Table 2) that provides reliable posterior estimates. We will reject all combinations of prior distributions that lead to cases of non-convergence in the posterior estimation or that tend to higher execution times or small effective posterior sample sizes. According to Kruschke, (2015), a posterior sample is recommended to be used for HDI+ROPE testing if the effective sample size is at least $10000$.
The second step includes a classification of tuning parameter combinations for the HDI+ROPE test and combination rule based on a sensitivity and specificity driven performance measure. The classification is restricted to prior distributions selected in step one. The performance of the implemented test setups (depending on ROPE and posterior CI definition, credibility level and combination rule) averaged over varying simulation scenarios (sample size of data, background rate, ADR rate, expected event-time) are measured in terms of the area under the curve (AUC) of the receiver operating characteristic graph with one threshold (Fawcett,, 2004; Lloyd,, 1998) based on equal numbers of ADR and no-ADR scenarios (and thereby equal numbers of true positives and negatives). The AUC is calculated using the R package ROCR (Sing et al.,, 2005).
In the third step, the BPgWSP test with the tuning parameters chosen in step two is investigated further. The influence of the sample scenario on the AUC performance is presented in terms of the AUC stratified with respect to sample size, ADR rate, and expected event-time. Finally, the robustness to miss-specification is explored by examining the AUC when the prior belief differs from the truth because the expected event-time is incorrectly specified.
3.5 Results of the simulation study
Per prior distributional choice (see Table 2) 8400 posterior estimations (resulting from 336 simulation scenarios divided by number of distribution choices times 100 repetitions) were conducted. In the following, the prior distributions are compared with respect to convergence issues, execution times and effective sample sizes.
Non-convergence in posterior estimation:
Only few posterior estimations failed due to non-convergence of the MCMC. This occurred for $3.6\%$ of the posterior estimations given a fix-gam-gam prior and for $10.2\%$ of the posterior estimations when a gam-gam-gam prior was used. No implications for convergence issues (in form of an error message and canceled execution of posterior estimation) came up when a fix-log-log or log-log-log prior was specified. As they seem to provide smooth estimation without the need to adjust the MCMC sampler, we suggest to use a fix-log-log or a log-log-log prior distribution.
Execution times:
Figure 5: Boxplots of running times split with respect to prior distributional choice. The graph is cut on the y-axis at 210 minutes (3h) although the boxplot for the gam-gam-gam prior reaches as maximum at about 45 hours.
A boxplot of the execution times for posterior estimation stratified with respect to the prior distribution is shown in Figure 5. All distribution choices lead to acceptable median execution times of 8 min for fix-log-log models, 13 min for log-log-log models, 10 min for fix-gam-gam models, and 40 min for gam-gam-gam models. While the maximal running times that occurred during the simulation stayed under about 1 h given a fix-log-log, 2 h given a fix-gam-gam, or 3 h given a log-log-log distribution, the gam-gam-gam prior produced execution times of up to 45 hours.
Effective sample sizes:
Figure 6: Boxplots of estimated effective sample size split with respect to prior distributional choice for parameter $\nu$ and $\gamma$.
In $73\%$ of the posterior estimations for $\nu$ and $55\%$ of the posterior estimations for $\gamma$ does the log-log-log model generate a posterior sample with an effective sample size $>10000$. The fix-log-log, fix-gam-gam and gam-gam-gam prior distributions provide only $\approx 25-28\%$ simulations for $\nu$ and $21-26\%$ for $\gamma$ fulfilling the rule of thumb (see Figure 6). This makes the log-log-log prior preferable compared to the alternative prior distributions.
The log-log-log distribution provides the best results in terms of convergence of the posterior estimation, adequate running times and sufficiently high effective posterior sample size in most cases. Consequently, we now proceed to evaluate the simulation results of the the log-log-log prior distribution.
Ranking of test specifications based on the AUC:
We rank all combinations of ROPE and posterior CI specifications, credibility levels and combination rules based on the AUC averaged over all sample scenarios given a correct prior belief. Table 4 shows the top ten test specifications in terms of AUC. The highest AUC value of $0.89$ is achieved with an $80\%$ Lognormal ETI as ROPE, an $80\%$ posterior HDI as posterior CI and option 2 for the combination rule (see Table 3).
rank ROPE posterior CI combination rule AUC
1 80% ETI 80% HDI option 2 0.890
2 80% ETI 80% ETI option 2 0.873
3 80% ETI 80% ETI option 3 0.838
4 80% ETI 80% HDI option 3 0.834
5 75% ETI 75% HDI option 3 0.823
6 75% ETI 75% ETI option 3 0.802
7 85% ETI 85% HDI option 2 0.693
8 85% ETI 85% ETI option 2 0.691
9 90% ETI 90% HDI option 1 0.653
10 85% ETI 85% HDI option 1 0.643
Table 4: Ranking of top ten BPgWSP test specifications according to AUC averaged over all sample scenarios given a correct prior belief.
Effect of simulation scenario on AUC performance:
The influence of the sample size, the proportion of ADR events and the expected event-time of the ADR (with correctly specified prior belief) is summarized in Table 5. The BPgWSP test works better with increasing sample size reaching an AUC of $98.9\%$ given a sample size of 5000. An increase in the proportion of ADR events seems to improve detectability of the ADR very slightly. The test performance is higher for ADRs occurring around the first or third quarter of the observation period with an AUC of $94\%$, but still provides an AUC of $78\%$ for ADRs occurring in the middle of the observation period.
grouping factors AUC
sample size $N$
$500$ 0.749
$3000$ 0.932
$5000$ 0.989
proportion of ADR events
$5\%$ of $N$ 0.881
$10\%$ of $N$ 0.899
expected time of ADR events
1st quarter of observation period 0.946
2nd quarter of observation period 0.782
3rd quarter of observation period 0.942
Table 5: Average AUC performance of the BPgWSP test (prior distribution: log-log-log, ROPE: $80\%$ Lognormal confidence interval, credibility interval: $80\%$ highest posterior density region, combination rule option 2) grouped by sample size, proportion of ADR events, and expected time of ADR events.
Robustness to miss-specification:
If the specification of the expected event-time is off by one quarter of the observation time, the AUC reduces to $78\%$ on average. If the a priori expected event-time is off by two quarters, the AUC performance drops to $47\%$. An AUC performance of $79\%$ is provided in case no ADR is suspected a priori.
For application of the BPgWSP test to a range of drug-AE pairs (in an automated manner), we recommend to set up the tuning parameters as follows: an $80\%$ Lognormal ETI for the ROPE, an $80\%$ posterior HDI for the posterior CI and option 2 for the combination rule. In the next section, the developed method will be applied to real data for demonstration purposes using the recommended tuning for the BPgWSP test.
type of prior belief AUC
correct specification 0.890
one quarter off 0.780
two quarters off 0.469
no ADR assumed 0.788
Table 6: AUC performance given a correct or incorrect prior belief of different types about the expected event-time.
4 Case study
The purpose of this case study is to illustrate the BPgWSP test derived in this work with a real dataset. Conclusions from this case study should not be used to inform clinical practice. The dataset was originally created from the Health Improvement Network (THIN) database (THIN,, 2024) to assess the association between the prescription of bisphosphonates and the occurrence of Carpal tunnel syndrome in menopausal Women (Carvajal et al.,, 2016). The THIN database contains primary care data from patients in the UK.
The BPgWSP test is applied to investigate the association between bisphosphonates and four AEs occurring in the population. In addition to the exposed cohort, the dataset contains patients not exposed to bisphosphonates (control group). A control group is not required to apply the BPgWSP test. Nevertheless, it provides us with the opportunity to explore the type one error, i.e. the test results for the control group.
The data preparation has been described in Sauzet et al., (2013). To facilitate the description of the data, we recall the descriptive statistics provided in Sauzet et al., (2013). All steps are conducted for the exposed and the control groups. We present the sample size and number of events as well as the mean event-time (and standard deviation). A histogram of event-times gives insights into the distribution over time. A smoothed non-parametric estimate of the hazard function provides information on the trend of risk for an AE to occur over time.
The BPgWSP test is applied with the recommended tuning parameter resulting from the simulation study (log-log-log priors, 80% Lognormal confidence interval for the ROPE, 80% HDI for the posterior CI, option 2 for the combination rule). Selectively, we compare our test outcomes to the test results obtained with the non-Bayesian approaches illustrated in Sauzet et al., (2013) and Sauzet and Cornelius, (2022).
4.1 Headache
Figure 7: Graphics for exploratory analysis of the occurrence of headache over 21 days. The graphic derived is based on Sauzet et al., (2013, Fig. 1).
Exploratory analysis:
The data is censored after 21 days. In this time, 21 among 19791 patients $(0.11\%)$ exposed to bisphosponates experienced headaches. The day of occurrence is on average $14\;(5.9)$. The histogram of event-times (Figure 7 a.1) shows that the number of events is slightly higher in the second half of the observation period. The non-parametric smooth hazard estimate (Figure 7 b.1) indicates a more pronounced increase of risk in the last seven days.
Within the control group, headache occurred in 27 among 39630 patients $(0.07\%)$. The events are distributed over the full observation window (see Figure 7 a.2). The hazard estimate is almost horizontal (see Figure 7 b.2).
prior belief about prior means test outcome test outcome
expected event-time $(\bar{\theta}_{0},\bar{\nu}_{0},\bar{\gamma}_{0})$ (Exposed) (Control)
none $(1,1,1)$ signal -
1st quarter of OP (5.25 days) $(2.5,0.5,1)$ signal -
2nd quarter of OP (10.5 days) $(3,3,5)$ signal -
3rd quarter of OP (15.75 days) $(18,5,1)$ signal -
Table 7: Test outcomes under various prior beliefs for headache based on exposed and control group.
Figure 8: Prior distribution and posterior sample histogram obtained from Bayesian model under the prior belief that headache is no ADR. The test results in row one of Table 7 are concluded from the alignment of the ROPE and HDIs and combination according to combination rule 2.
Signal detection test:
The prior specifications and results of the BPgWSP test are shown in Table 7. The signal detection tool is applied to the data censored at 21 days under the prior beliefs that the symptom occurrence is expected at the first, second or third quarter of the observation period, as well as a prior belief of no association. Independent of the prior specification including the assumption of no association, a signal is consistently raised for the exposed group and no signal for the control group. This is in line with the results of the non-Bayesian approach.
The intermediate shape parameters’ HDI+ROPE test outcomes are illustrated in Figure 8 for the exposed and control group under the prior belief of no association. In the case of the exposed cohort, both single parameter tests reject the null-hypothesis as the HDI does not lie within the ROPE) and combination scheme option 2 concludes that a signal is raised. In contrast, the control group’s HDI+ROPE test for $\gamma$ returns a rejection of the null-hypothesis while the HDI+ROPE test for $\nu$ returns an acceptance. According to the combination scheme option 2, no signal is raised.
4.2 Musculoskeletal pain
Figure 9: Graphics for exploratory analysis of the occurrence of musculoskeletal pain over one year. The graphic derived is based on Sauzet et al., (2013, Fig. 2).
Exploratory analysis:
The data is censored after one year in which 207 of 19777 ($1\%$) exposed patients reported musculoskeletal pain. The average event-time to occurrence is 160 days $(99)$. The histogram of event-times (Figure 9 a.1) shows a slightly higher density in the first half of the observation period. The smoothed hazard estimate (Figure 9 b.1) indicates a slight increase in risk during the first four months followed by a decrease.
In the control group musculoskeletal pain was reported 207 times from 39634 patients ($0.5\%$). According to the histogram (Figure 9 a.2) slightly more cases are reported in the first and third month but evenly distributed for the rest of the time. The hazard estimate (Figure 9 b.2) looks almost constant.
prior belief about prior means test outcome test outcome
expected event-time $(\bar{\theta}_{0},\bar{\nu}_{0},\bar{\gamma}_{0})$ (Exposed) (Control)
none $(1,1,1)$ signal -
1st quarter of OP (91 days) $(1,0.207,1)$ signal -
2nd quarter of OP (183 days) $(20,5.5,14)$ signal -
3rd quarter of OP (274 days) $(300,4,1)$
Table 8: Test outcomes under various prior beliefs for musculoskeletal pain based on exposed and control group.
Signal detection test:
The results of the BPgWSP test applied to the data censored at 365 days under the prior beliefs of symptom occurrence at the first, second or third quarter of the observation period or of no association are shown in Table 8. A signal is raised for the exposed group under all priors except a prior belief around the third quarter (day 274). For the control group, no signal is raised regardless of the prior choice.
In the case of musculoskeletal pain, the test results depend on the prior specification. Here we have a long observation period and a small number of cases which potentially decreases the performance of the BPgWSP test. Given that most cases occur at the beginning of the observation period the prior specification for the last observation quarter does not lead to a signal. This presumably occurs because data and prior belief provide contrasting information.
4.3 Alopecia
Figure 10: Graphics for exploratory analysis of the occurrence of alopecia over three years. The graphic derived is based on Sauzet et al., (2013, Fig. 3).
Exploratory analysis:
Over three years there have been 144 reports of alopecia in the group of 19812 patients (0.73%) exposed to bisphosphonates. The mean event-time is on day 424 (285). According to the histogram (Figure 10 1.a) events tend to take place slightly more often between the fifth and 13th month which roughly corresponds to the period around the first or second quarter of the observation period. Regarding the estimated hazard function the risk rises and falls within the first two years and stabilizes after that.
There are 180 alopecia cases among 39634 control patients ($0.45\%$). The event-times are slightly more dense in the first half of the observation window. The estimated hazard (Figure 10 2.b) has an almost flat trend.
Signal detection test:
prior belief about prior means test outcome test outcome
expected event-time $(\bar{\theta}_{0},\bar{\nu}_{0},\bar{\gamma}_{0})$ (Exposed) (Control)
none $(1,1,1)$ signal -
1st quarter of OP (274 days) $(200,0.63,1)$ signal -
2nd quarter of OP (548 days) $(30,3.2,10)$ signal -
3rd quarter of OP (821 days) $(700,14,4)$
Table 9: Test results under various prior beliefs for alopecia based on exposed and control group.
Table 9 shows the BPgWSP test results for alopecia. Under the prior assumptions of no association, ADR occurrence in the beginning or in the middle of the observation period the test identifies a temporal pattern for the exposed group and none for the control group. For both groups, no signal is raised under the prior belief that alopecia risk increases at the end of the observation period.
Alopecia is a rare event. In the non-Bayesian approach in Sauzet et al., (2013) an ambiguous signal was given if the data was censored after one year only whereas here, we get a consistent signal if the prior belief does not expect cases to occur toward the three-year mark. For the control group, no signal is raised indicating that the signal raised for cases is probably not due to noise.
4.4 Carpal tunnel syndrome
Figure 11: Graphics for exploratory analysis of the occurrence of carpal tunnel syndrome over three years. The graphic derived is based on Sauzet et al., (2013, Fig. 4).
Exploratory Analysis:
Within the observation period of three years, 192 among 19817 exposed patients ($0.97\%$) experienced carpal tunnel syndrome. The day of occurrence is on average $429\;(5.9)$ which is approximately one year after drug prescription. The histogram of event-times (Figure 11 a.1) reveals that events very slightly decrease over the scope of the observation period. The hazard estimate (Figure 11 b.1) indicates a slightly higher constant risk in the first 18 months and a decrease after that.
Within the control group, carpal tunnel syndrome occurred in 244 among 39389 patients ($0.62\%$). The distribution of events looks similar to the exposed group’s plot (see Figure 7 a.2). The hazard estimate shows an incline of risk in the first year, a rather constant trend in the middle and again an increase within the last 5 months. (see Figure 11 b.2).
Signal detection test:
prior belief about prior means test outcome test outcome
expected event-time $(\bar{\theta}_{0},\bar{\nu}_{0},\bar{\gamma}_{0})$ (Exposed) (Control)
none $(1,1,1)$
1st quarter of OP (274 days) $(200,0.63,1)$ signal signal
2nd quarter of OP (548 days) $(30,3.2,10)$ signal signal
3rd quarter of OP (821 days) $(700,14,4)$
Table 10: Test outcomes under various prior beliefs for carpal tunnel syndrome based on exposed and control group.
The BPgWSP test results for carpal tunnel syndrome are presented in Table 10. In two out of four prior specifications, namely when the prior belief is that the event is to be expected around the first quarter or middle of the observation period, a signal is raised for both control and exposed group. In the two other cases, a signal is raised for neither group.
The ambiguous test results occurring here and in Sauzet et al., (2013) where no signal was raised for the exposed group and a signal was raised for the control group when censored after 637 days are probably returned due to the confounding effect of the age of the participants for which the carpal tunnel syndrome is a common condition.
5 Discussion
In this work we developed the BPgWSP test that raises a signal if there is an indication of non-constant hazard over time for an AE after drug intake incorporating prior knowledge about the time of occurrence. The defined test was tuned based on a simulation study which led to a one-fits-all approach. Consequently, we applied the test with the recommended tuning to real data.
The underlying assumption of the BPgWSP test is that the hazard should be constant if there is no drug-AE association. This assumption also underlies the work of Cornelius et al., (2012) and the extension by Sauzet and Cornelius, (2022) in the frequentist framework. While this heuristic assumption can be refuted due to various confounding factors, we argue that it is a simplified model is good enough for flagging possible associations because finding causal association is not the aim of the procedure.
The simulation study enabled an initial assessment of the test performance and the conclusion of a recommended BPgWSP test tuning for application in an automated one-fits-all manner to a multitude of drug-AE combinations. Arguably, the amount of combinations of simulation scenarios and prior specifications considered within the frame of the simulation study is too limited to draw general conclusions on tuning parameters and only give advise for tuning choices. For instance only one censoring time (365) was considered for data generation and only one standard deviation (10) was considered for the prior belief specification such that insights on the expected performance in case of alternative specifications cannot be drawn. This was due to the practicality of presenting the methods. In general however, we recommend to tune the BPgWSP test with an additional simulation study adjusted for the circumstances of interest. To do so, we provide the R package BWSPsignal available on https://github.com/julia-dyck/BWSPsignal such that new values to specify the data generating characteristics (see sec. 3.1), model specifications (see sec. 3.2) and tuning parameters for the HDI+ROPE test setup (see sec. 3.3) can be easily evaluated.
In the case study, we applied the BPgWSP test to real data on bisphosphonates with adverse events known to be ADRs caused by the drug (headache, musculoskeletal pain) and adverse events not identified as ADR (alopecia, carpal tunnel syndrome). The BPgWSP test tuned based on the simulation results was able to raise a signal under varying prior beliefs for the known ADRs. It returned ambiguous test results for adverse events not categorized as ADRs under some prior belief specifications. The performance of the BPgWSP test depends on whether the observation period encompasses the expected event-time and on the rareness of the adverse event. This is an additional argument for the need of more differentiated and context specific tuning of the BPgWSP test. Simulations representing the data scenarios better could improve the test performance in the case study applications.
The data used for the case study is a historical dataset derived from the THIN database (THIN,, 2024). It would be desirable to test the newly developed methods with a larger set of drug-AE pairs. However, EHR data are not easily available for method development due todata privacy regulations and costs. There are attempts to create synthetic, yet realistic datasets. An example is the SYNTHEA database (Walonoski et al.,, 2017). Even though the SYNTHEA database is based on a complex data-generating process including modules for a set of health conditions and medications that occur with probabilities and prevalences reflecting reality, the generating process does not include AEs. We think that further development of SYNTHEA and similarly structured databases needs to be supported even more extensively to facilitate the development of statistical methods for pharmacoepidemiology and other fields.
6 Conclusion
Signal detection is an important component of pharmacovigilance to monitor drugs post market release. To improve further development of statistical signal detection methods, we combine two characteristics of existing approaches in the literature, the use of longitudinal data from EHRs and the incorporation of prior knowledge about the drug under investigation, in one test.
The Bayesian WSP test is based on the constant-hazard assumption and consists of fitting a Bayesian PgW model with a prior distribution for the PgW parameters reflecting the prior belief about the drug-AE relationship, as well as a tuned HDI+ROPE test based on the posterior shape parameters. We showed how to tune the parameters specif using a simulation study, and illustrated it on data about bisphosphonate prescriptions.
Data availability statement
Funding
No funding was received to perform this study.
Conflicts of interest
The authors declare no conflict of interest.

### v3

@@
 Abstract
-We develop a Bayesian Power generalized Weibull shape parameter (PgWSP) test as statistical method for signal detection of possible drug-adverse event associations using electronic health records for pharmacovigilance.
+We develop the Bayesian Power generalized Weibull shape parameter (BPgWSP) test as statistical method for signal detection of possible drug-adverse event associations using electronic health records for pharmacovigilance.
 The Bayesian approach allows the incorporation of prior knowledge about the likely time of occurrence along time-to-event data.
 The test is based on the shape parameters of the Power generalized Weibull (PgW) distribution.
-When both shape parameters are equal to one, the PgW distribution reduces to an exponential distribution, i.e.
-a constant hazard function.
-This is interpreted as no temporal association between drug and adverse event.
-The Bayesian PgWSP test involves comparing a region of practical equivalence (ROPE) around one reflecting the null hypothesis with estimated credibility intervals reflecting the posterior means of the shape parameters.
-The decision to raise a signal is based on the outcomes of the ROPE test and the selected combination rule for these outcomes.
-The development of the test requires a simulation study for tuning of the ROPE and credibility intervals to optimize specificity and sensitivity of the test.
-Samples are generated under various conditions, including differences in sample size, prevalence of adverse drug reactions (ADRs), and the proportion of adverse events.
+When both shape parameters are equal to one, the PgW distribution reduces to an exponential distribution, yielding a constant hazard function.
+This is interpreted as no temporal association between drug and adverse event (AE).
+The BPgWSP test involves comparing a region of practical equivalence (ROPE) around one reflecting the null hypothesis with estimated credibility intervals (CI) reflecting the posterior means of the shape parameters.
+The decision to raise a signal is based on the CI+ROPE tests and the selected combination rule for these outcomes.
+The test development requires a simulation study for tuning of the ROPE and CIs to optimize specificity and sensitivity of the test.
+Samples are generated under various conditions, including differences in sample size, prevalence of adverse drug reactions (ADRs), and the proportion of AEs.
 We explore prior assumptions reflecting the belief in the presence or absence of ADRs at different points in the observation period.
-Various types of ROPE, credibility intervals, and combination rules are assessed and optimal tuning parameters are identified based on the area under the curve.
-The tuned Bayesian PgWSP test is illustrated in a case study in which the time-dependent correlation between the intake of bisphosphonates and four adverse events is investigated.
+Various types of ROPE, CIs, and combination rules are assessed, and optimal tuning parameters are identified based on the area under the curve.
+The tuned BPgWSP test is illustrated in a case study in which the time-dependent correlation between the intake of bisphosphonates and four AEs is investigated.
+Julia Dyck [^1] Department of Business Administration and Economics, Bielefeld University, Universitätsstraße 25, Bielefeld, 33615, Germany.***Corresponding author: j.dyck@uni-bielefeld.de and Odile Sauzet11footnotemark: 1222School of Public Health, Bielefeld University, Universitätsstraße 25, Bielefeld, 33615, Germany.
 Key words:
@@
 1 Introduction
-Making assumptions about the generating process of an outcome in the development of a test of association can improve the performance of this test in terms of power, specificity, and sensitivity by enabling the use of more statistical information included in the data.
-In the field of pharmacovigilance (surveillance of occurrence of adverse events (AEs) related to prescribed drugs (Meyboom et al.,, 1997), this approach has been used by Cornelius and colleagues (Cornelius et al.,, 2012; Sauzet and Cornelius,, 2022) to develop a family of signal detection tests (test of association between the occurrence of an AE and a prescribed drug) which is based on an assumption about time dependence on drug prescription.
-More specific, Cornelius et al., (2012) proposed tests aiming for detection of a non-constant hazard function by modeling the time to occurrence of event from the time of prescription to flag a possible drug-event association.
-However, this is not the only type of information about a generating process that can be incorporated into a test.
+In the field of pharmacovigilance, signal detection methods are used to identify new potential adverse drug reactions1 (ADRs) among adverse events 2 (AEs) occurring in the population by testing for an association between the occurrence of an AE and a prescribed drug.3 Signal detection is important to keep a drug‘s harm profile updated and can result in adjustments of the prescription labelling or even a recall of the product from the market.
+4, 5, 6
+With the collection of longitudinal electronic health records7, 8 (EHRs) (regularly collected data about diagnoses, treatments, and medication prescribed by physicians, as well as AEs), statistical methods to exploit them have emerged.
+These include disproportionality approaches adjusted for longitudinal data comparing rates of occurrence between a single drug and the rest of the dataset 9, 10, temporal pattern analysis 5, exposure models 11 and more (see for example the review by Coste et al.
+12).
+Sauzet and Cornelius13, 6 developed a family of signal detection tests based on an assumption about time dependence on the drug prescription.
+More specific, their tests aim for detection of a non-constant hazard function (depicting the risk over time) by modelling right-censored time-to-event data starting at drug prescription to flag a possible drug-event association.
+Making assumptions about the generating process of an outcome (more specifically the hazard function) in the development of a test of association can improve the performance of this test in terms of power, specificity, and sensitivity by enabling the use of more statistical information included in the data.
+However, this is not the only type of information that can be incorporated into a test.
 Prior knowledge regarding the distribution parameters of the generating process (e.g.
-mean time of occurrence of an event) can be incorporated in the development of a test using a Bayesian approach.
-The aim of this article is to combine both approaches by developing a Bayesian test to detect non-constant hazard functions for a time-dependent process incorporating prior knowledge on the distribution of the time to the occurrence of events.
-While the principle is rather general, we present this test in the context of pharmacovigilance which we briefly describe below, and review some existing methods of signal detection that assume a temporal relationship or use a Bayesian approach.
-After release of a drug on the market, pharmacovigilance monitors the occurrences and changes of known adverse drug reactions (ADRs) as well as aims to detect new ADRs in the population.
-An ADR is an unintended, harmful response to a medicine (EMA,, 2024).
-The investigation of a harm profile of a drug consists of multiple steps: screening the range of AEs that occur along prescribed drugs, performing a signal detection test to investigate the association for a drug-AE pair and, in case a signal is raised, deeper investigation of any causal relationship between the two which involves a complex epidemiological analysis (Meyboom et al.,, 1997).
-Longitudinal electronic health records (EHRs) gathered by health insurance companies and health regulatory organizations such as the Sentinel system established by the FDA (Food and Drug Administration,, 2022) provide regularly collected longitudinal data about diagnoses, treatments, and medication prescribed by physicians, as well as AEs (Suling and Pigeot,, 2012; Sturkenboom and Schink,, 2021).
-With the collection of EHR data, statistical methods to exploit them have emerged.
-These include among others extensions of disproportionality approaches comparing rates of occurrence between a single drug and the rest of the dataset (Zorych et al.,, 2013) for application to longitudinal data (Schuemie,, 2011), temporal pattern analysis (Norén et al.,, 2010), exposure models (Dijkstra et al.,, 2024) and more (see for example the review by Coste et al., (2023)).
-As mentioned above Cornelius et al., (2012) and Sauzet and Cornelius, (2022) considered the temporal dynamics of the risk to experience a specific AE in their family of Weibull shape parameter (WSP) family of tests.
-For this a Weibull type distribution parametric hazard function is fitted to the data.
-A signal is raised following the rejection of a null-hypothesis assuming a constant hazard which means that a sufficiently large change in risk over time can be interpreted as a signal.
-Unlike disproportionality methods, no group comparisons are performed.
-Bayesian statistical methods for signal detection have also been developed.
-But methods like the Empirical Bayes Gamma Mixture/ Gamma-Poisson shrinker or the Bayesian confidence propagation neural network are Bayesian extensions of disproportionality measures used to obtain signals from spontaneous reporting databases (Bansal et al.,, 2024) which do not consider the time of the event occurrence.
-The objective of this work is to combine an assumption of temporal process with the incorporation of prior knowledge about distribution parameters of this process in the development of Bayesian Weibull shape parameter tests.
-To do so, we will generalize the proposed power generalized Weibull shape parameter (PgWSP) test by Sauzet and Cornelius, (2022) into a Bayesian framework to develop and tune an HDI+ROPE type test of signal detection.
-The remaining article is structured as follows.
-In section 2 the theoretical concept of the Bayesian PgWSP (BPgWSP) test is presented.
-In section 3, we conduct a simulation study to tune the parameters defining the BPgWSP test in a chosen framework of simulation scenarios.
-Finally, in section 4, we apply our tuned BPgWSP test to data on women prescribed with bisphosphonates from the The Health Improvement Network (THIN) (THIN,, 2024) database in a case study.
+mean time of occurrence of an event) might be incorporated into the development of a test using a Bayesian approach.
+Bayesian statistical methods like the Empirical Bayes Gamma Mixture, also called Gamma-Poisson shrinker, or the Bayesian confidence propagation neural network are extensions of disproportionality measures used to obtain signals from spontaneous reporting databases.14, 15 These, in turn, do not consider the time of the event occurrence.
+This article aims to combine the assumption of a temporal process and the incorporation of prior knowledge about distribution parameters of this process.
+To do so, we will generalize the power generalized Weibull shape parameter (pWSP) test proposed by Sauzet and Cornelius 6 into a Bayesian framework to develop and tune an HDI+ROPE16 type test (a Bayesian equivalent of a hypothesis test) of signal detection which we call the Bayesian Power generalized Weibull shape parameter (BPgWSP) test.
 2 Methodological concept for the Bayesian signal detection test
@@
 The principle of non-constant hazard signal detection tests is built on the assumption that the risk of an AE is time-dependent relative to the prescription of the drug under investigation.
+The method is conceptualized for right-censored time-to-event data.
+Each observation $i$ consists of the time $t_{i}\in(0,c)$ with $c$ being the censoring time and the binary status $d_{i}\in\{0,1\}$ indicating whether the AE occurred at $t_{i}$.
 The parametric distribution is defined by a parameter vector $\Theta$.
-We make use of a subset of parameters $\mathcal{N}\subset\Theta$ on which to define the null-hypothesis of constant hazard and develop the Bayesian signal detection test based on an updated distribution for $\mathcal{N}$ using the HDI+ROPE approach (see sec.
+We define the null-hypothesis of constant hazard based on a subset of parameters $\mathcal{N}\subset\Theta$ and develop the Bayesian signal detection test (see Figure 1) based on an updated distribution for $\mathcal{N}$ using the HDI+ROPE approach (see sec.
 2.2).
-For this we will perform the three steps below:
+For development, we will perform the three steps below:
   1. 1.
@@
   2. 2.
-    describe the BPgWSP test using the HDI+ROPE approach,
+    describe the BPgWSP test using the HDI+ROPE approach based on a Bayesian estimation,
   3. 3.
     set the prior specification and HDI+ROPE tuning parameters for the BPgWSP test by performing a simulation study (see sec. 3).
-2.1 Power generalized Weibull distribution:
-The survival function of the PgW distribution is given by:
+2.1 Power generalized Weibull distribution
+The survival function of the PgW distribution17, 18 is
 $S(t)=\text{exp}\left\{1-\left[1+\left(\frac{t}{\theta}\right)^{\nu}\right]^{\frac{1}{\gamma}}\right\}.$ (1)
-The hazard function is given by:
-$h(t)=\frac{\nu}{\gamma\theta^{\nu}}\cdot t^{\nu-1}\cdot\left[1+\left(\frac{t}{\theta}\right)^{\nu}\right]^{\frac{1}{\gamma}-1},$ (2)
-(see Bagdonavicius and Nikulin, (2001, p.8); Nikulin et al., (2016, p.8)).
-The PgW distribution has three parameters $\Theta_{PgW}=\{\theta,\nu,\gamma\}$ two of which are shape parameters such that $\mathcal{N}_{PgW}=\{\nu,\gamma\}$.
+The hazard function17, 18 is
+$h(t)=\frac{\nu}{\gamma\theta^{\nu}}\cdot t^{\nu-1}\cdot\left[1+\left(\frac{t}{\theta}\right)^{\nu}\right]^{\frac{1}{\gamma}-1}.$ (2)
+The PgW distribution has three parameters $\Theta_{PgW}=\{\theta,\nu,\gamma\}$ - two of which are shape parameters such that $\mathcal{N}_{PgW}=\{\nu,\gamma\}$.
 Figure 2: Hazard functions obtained from power generalized Weibull distributions with a) constant ($\theta=2,\nu=1,\gamma=1$) , b) increasing ($\theta=8,\nu=3,\gamma=1$), c) decreasing ($\theta=1.5,\nu=0.6,\gamma=1$), d) unimodal ($\theta=0.5,\nu=5,\gamma=10$), e) bathtub ($\theta=30,\nu=0.9,\gamma=0.2$) form.
-The two shape parameters allow for a range of hazard function forms including constant, monotoneously in- or decreasing and unimodal or bathtub shaped (see Figure 2).
-A constant hazard is given if and only if both shape parameters are equal to one so that the null-hypothesis of constant hazard can be formulated as:
+The two shape parameters allow for a range of hazard function forms including constant, monotoneously in- or decreasing, and unimodal or bathtub shaped (see Figure 2).
+A constant hazard is given if and only if both shape parameters are equal to one, so that the null-hypothesis of constant hazard can be formulated as
 $H_{0}:\nu=1\text{ and }\gamma=1.$ (3)
 2.2 Development of Bayesian signal detection test using the HDI+ROPE approach
-Prior and posterior distribution
-The prior distribution of $\Theta$ should reflect whether the AE of interest is suspected to be an ADR (non-constant hazard) and if so, when it is expected to occur (specifying $\mathbf{E}[t]$) a priori by medical experts.
-This can be translated into the corresponding expected hazard form (comp.
-Figure 2) parameterized in terms of prior mean values for the parameters in $\Theta$.
-The prior distribution of the scale parameter reflects the AE background rate.
-The priors of the shape parameters $\mathcal{N}\subset\Theta$ determine the constancy or non-constancy and general trend of the hazard over time.
-Based on the specified prior means (and a chosen standard deviation) as well as a distributional choice (eg.
-a Lognormal distribution for each parameter in $\Theta$), the prior distribution $p_{0}(\Theta)$ will be updated with the data available.
+The prior distribution of $\Theta$ should reflect whether the AE of interest is suspected to be an ADR (by choosing a parameter combination yielding a non-constant hazard) and if so, when it is expected to occur a priori by medical experts (by specifying parameters $\Theta$ that lead to the a priori intended $\mathbf{E}[t]$).
+The scale parameter reflects the AE background rate.
+The shape parameters $\mathcal{N}\subset\Theta$ determine the constancy or non-constancy and general trend of the hazard over time.
+For practicality, univariate prior distributions for each parameter out of $\Theta$ are specified.
+Concerning implementation, not $\mathbf{E}[t]$ but the prior means of all parameters are set, followed by checking whether the hazard form and $\mathbf{E}[t]$ resulting from the chosen prior means reflect the prior belief to a sufficient extent in a trial-and-error manner.
+Based on the specified prior means (and a chosen standard deviation) as well as a distributional choice (e.g.
+the product of univariate lognormal distributions for each parameter in $\Theta$), the prior distribution $p_{0}(\Theta)$ will be updated yielding a posterior distribution.
 For time-to-event data the posterior distribution $p_{1}(\Theta|t)$ is proportional to
 $p_{1}(\Theta|t)\propto p_{0}(\Theta)\cdot\mathcal{L}(t|\Theta)$ (4)
-where
+with
 $\mathcal{L}(t|\Theta)=\prod_{i=1}^{N}S(t_{i})^{1-d_{i}}\cdot f(t_{i})^{d_{i}}$ (5)
-with $\mathcal{L}(t|\Theta)$ being the likelihood function in the time-to-event framework such that all subjects $i$ not having an AE during the observation time (denoted by $d_{i}=0$) are considered in terms of the survival function $S(t_{i})$ and all subjects $i$ experiencing an AE (denoted by $d_{i}=1)$ are taken into account in terms of the density function $f(t_{i})=S(t_{i})\cdot h(t_{i})$.
-HDI+ROPE method
-Based on the posterior distribution of $\mathcal{N}$, the HDI+ROPE test (Kruschke,, 2015, 2018) applied to each shape parameter and combination of all single shape parameter test results provide two possible outcomes: either we lean towards the null-hypothesis of constant hazard or the alternative hypothesis of a non-constant hazard.
+being the likelihood function in the time-to-event framework such that all subjects $i$ not having an AE during the observation time (denoted by $d_{i}=0$) are considered in terms of the survival function $S(t_{i})$ and all $i$ experiencing an AE (denoted by $d_{i}=1)$ are taken into account in terms of the density function $f(t_{i})=S(t_{i})\cdot h(t_{i})$.
+Based on the posterior distribution of $\mathcal{N}$, the HDI+ROPE test 19, 16 is applied to each shape parameter and combined, leading to two possible outcomes: either we lean towards the null-hypothesis of constant hazard or the alternative hypothesis of a non-constant hazard.
 The test setup depends on
-  • the region of practical equivalence (ROPE) to the null value, in our case one, typically defined as an interval around the null value and
-  • a credibility interval (CI) based on posterior distribution parameters in $\mathcal{N}$, here either a highest density interval (HDI) as proposed by Kruschke, (2015) or an equal tailed interval (ETI).
+  • the region of practical equivalence (ROPE) to the null value (in our case, one), typically defined as an interval around the null value and
+  • a credibility interval (CI) based on posterior distribution parameters in $\mathcal{N}$, here either a highest density interval (HDI) as proposed by Kruschke19 or an equal-tailed interval (ETI).
 See sec.
 3.3 for the formal definition of the ETI and HDI.
-Figure 3: HDI+ROPE decision rules based on graphic from Kruschke, (2018, Fig. 1).
-In our setting, the HDI+ROPE test is applied to each parameter $\nu\in\mathcal{N}$ and provides one of the following interim results (compare to Figure 3):
+Figure 3: HDI+ROPE decision rules based on graphic from Kruschke, Fig. 1.16
+In our setting, the HDI+ROPE test is applied to each parameter $\nu\in\mathcal{N}$ and provides one of the following interim results:
   • If the CI lies completely within the ROPE, then the single parameter null-hypothesis
     $H_{0}(\nu):\nu=1$
-    is accepted,
-  • if ROPE and CI do not intersect at all, then the single parameter null-hypothesis is rejected,
-  • if ROPE and CI partly overlap, then no decision is made (undecided).
-Given the single parameter test outputs, the combined test result is concluded according to a combination rule, for instance the condition directly implied by Equation 3:
+    is accepted (Figure 3b-c),
+  • if ROPE and CI do not intersect at all, then the single parameter null-hypothesis is rejected (Figure 3a),
+  • if ROPE and CI partly overlap, no decision is made (Figure 3d-f).
+Given the single parameter test outputs, the combined test result is concluded according to a combination rule, for instance, the condition implied by Equation 3:
   • If the HDI+ROPE accepts the null-hypothesis for all parameters in $\mathcal{N}$, then accept the null-hypothesis of constant hazard.
   • Else, reject the null-hypothesis of constant hazard and raise a signal.
-To complete the test development, we need to make decisions about the distributional choice for the prior, the position and width of the ROPE and the CI of the BPgWSP test and the combination rule.
-This will be done in the following section.
+To complete the test development, we need to make decisions about the prior distributional choice, the position and width of the ROPE and the CI of the BPgWSP test and the combination rule.
 3 Simulation study for test development and tuning
-A method to finalize the BPgWSP test theoretically described in section 2 is to perform a simulation study.
-By applying a range of prior distributions, choices for position and width of the ROPE, position and width of the CI and the combination rules, we make decisions about the optimal tuning parameters for the BPgWSP test such that a performance parameter based on sensitivity and specificity of the test is maximal across a range of sample scenarios.
+A method to finalize the BPgWSP test, theoretically described in sec.
+2, is to perform a simulation study.
+By applying a range of prior distributions, choices for position and width of the ROPE, position and width of the CI, and combination rules, we make decisions about the optimal tuning parameters for the BPgWSP test such that a performance parameter based on sensitivity and specificity of the test is maximal across a range of sample scenarios.
+The frequentist pWSP test approach developed by Sauzet and Cornelius 6 with credibility levels to be tuned is considered in the simulations for comparison reasons.
 3.1 Data generating process
-The set of simulation scenarios is determined by the
-  • censoring time (365 days) which sets the end of the observation period,
+The sample scenarios are determined by the
+  • censoring time (365 days) setting the end of the observation period (OP),
   • sample size $N$ (500, 3000, 5000),
@@
   • ADR rate which is the proportion of the background rate (0, 0.5, 1),
-  • expected event-time of the ADR if the ADR rate is above zero (1st, 2nd, 3rd quarter of observation period),
+  • expected event-time of the ADR if the ADR rate is above zero (1st, 2nd, 3rd quarter of OP),
   • relative standard deviation of the ADR event-time (0.05).
-The background observations are sampled using a Binomial distribution with a probability equal to the specified background rate
+The background observations are sampled from a binomial distribution with a probability equal to the specified background rate
 $n^{br}\sim Bin(n=N,p=br).$
-Corresponding event-times are sampled using a Uniform distribution over the study period ranging from day one to day $365$.
+Corresponding event-times are sampled from a uniform distribution over the OP ranging from day one to $365$.
 $t^{br}$ $\sim Unif(1,365)$ (6)
-The number of events induced by the drug intake is sampled from a Binomial distribution with a probability equal to the background rate times the ADR rate.
+The number of events induced by the drug is sampled from a binomial distribution with the probability being the background rate times ADR rate.
 $n^{ADR}\sim Bin(n=N,p=br\cdot adr)$ (7)
-The event-times for the ADR cases are generated from a Normal distribution with an expected event-time $\mathrm{E}[t]$ at the first quarter, half, and third quarter of the study period and a standard deviation of $0.05$ times the censoring time at 365 days.
-$t^{ADR}$ $\sim N(\mathrm{E}[t],0.05\cdot 365)$ (8)
-The sample is filled up to size $N$ with censored observations. They have the AE status zero and time point $365$ equaling the end of the study period.
-3.2 Prior specifications for the Bayesian model
-Figure 4: Hazard functions obtained under the prior belief that the expected event-time is at the first, second, or third quarter of the observation period (left) or that the AE of interest is no ADR (right).
-prior belief about prior means
-expected event-time $(\bar{\theta}_{0},\bar{\nu}_{0},\bar{\gamma}_{0})$
-none $(1,1,1)$
-1st quarter of OP (91 days) $(1,0.207,1)$
-2nd quarter of OP (183 days) $(20,5.5,14)$
-3rd quarter of OP (274 days) $(300,4,1)$
-Table 1: Mean of prior distribution under implemented prior beliefs about expected event-time.
+Event-times for the ADR cases are generated from a normal distribution with an expected event-time $\mathrm{E}[t]$ at the first, second, and third quarter of the OP and a standard deviation of $0.05$ times the censoring time at 365 days.
+$t^{ADR}$ $\sim\mathcal{N}(\mathrm{E}[t],(0.05\cdot 365)^{2})\$(8)
+The sample is filled up to size $N$ with censored observations. They have the AE status zero and time point $365$ (end of the OP).
+3.2 Prior specifications and model estimation
+Figure 4: Hazard functions obtained under the prior belief that the expected event-time is at the first, second, or third quarter of the OP (left) or that the AE of interest is no ADR (right).
 Four different prior beliefs are implemented.
-An ADR is expected to occur with higher risk either around the first, second or third quarter of the observation period or not expected at all.
+An ADR is expected to occur with higher risk either around the first, second, or third quarter of the OP or not expected at all.
 These assumptions are represented by hazard functions of the distribution shown in Figure 4 based on chosen prior mean values for the scale $\theta$, and shapes $\nu$ and $\gamma$ listed in Table 1.
-The standard deviation of the parameters’ prior distributions is set to 10.
-scale $\theta$ shape $\nu$ powershape $\gamma$ abbreviation
-fixed to prior mean Gamma Gamma fix-gam-gam
-Gamma Gamma Gamma gam-gam-gam
-fixed to prior mean Lognormal Lognormal fix-log-log
-Lognormal Lognormal Lognormal log-log-log
-Table 2: Prior distributional assumptions for the scale $\theta$, and shapes $\nu$ and $\gamma$ of the PgW distribution.
-For practical reasons, scale and shape parameters are assumed to be independent.
-By definition all parameters of the PgW distribution are positive.
-Four prior distributional choices matching these conditions are listed in Table 2.
-All parameters’ univariate priors follow either the Gamma or Lognormal distribution.
-In two of the four prior distribution choices, the scale parameter of the posterior remains equal to the prior (i.e.
-will not be updated).
-For instance, the PgW model with log-log-log prior is be specified as follows:
-tsimilar-$t\sim$ $PgW(\theta,\nu,\gamma)$ (9)
-$\theta_{0}\sim Lognormal(\mu_{0}^{\theta},\sigma_{0}^{\theta})$ (10)
-$\nu_{0}\sim Lognormal(\mu_{0}^{\nu},\sigma_{0}^{\nu})$ (11)
-$\gamma_{0}\sim Lognormal(\mu_{0}^{\gamma},\sigma_{0}^{\gamma})$ (12)
-where the prior meta parameters of the Lognormal distribution are the logarithm of the location $\mu$ and the logarithm of the scale $\sigma$ which can be calculated from the prior mean (specified in Table 1 and standard deviation (10) as
-$\mu_{0}^{\theta}$ $=\log(\bar{\theta}_{0})-\log\left(\frac{sd(\theta_{0})^{2}}{\bar{\theta}_{0}^{2}}+1\right)$ (13)
-$\sigma_{0}^{\theta}$ $=\sqrt{\log(\frac{sd(\theta_{0})^{2}}{\bar{\theta}_{0}^{2}}+1)}$ (14)
-for $\theta$ and analogously for $\nu$ and $\gamma$.
-A posterior estimate is obtained by Markov chain Monte Carlo (MCMC) sampling.
-We draw a posterior sample of size 40000 obtained from four chains of length 10000 after a burn-in phase of 1000 draws each.
-Each combination of sample scenario and prior specification leads to a total of 336 different simulation settings – 288 with a positive ADR rate and 48 controls to match in with only background events for the simulation study.
+The standard deviation of all parameters is set to 10.
+See the supplementary material for details on the implementation.
+prior belief about a priori prior means
+expected event-time $E[t]$ $(\bar{\theta}_{0},\bar{\nu}_{0},\bar{\gamma}_{0})$
+none - $(1,1,1)$
+1st quarter of OP 91 days $(1,0.207,1)$
+2nd quarter of OP 183 days $(20,5.5,14)$
+3rd quarter of OP 274 days $(300,4,1)$
+Table 1: Mean of prior distribution under implemented prior beliefs about expected event-time. Prior standard deviation is set to 10 for all parameters in all prior belief cases.
+To obtain the posterior estimate the No-U-Turn sampler is used 20.
+We draw a posterior sample of size 40 00040\,000 obtained from four chains of length 10 00010\,000 after a burn-in phase of 1 0001\,000 draws per estimation.
+In practice, model diagnostics should be used to check convergence and sampling efficiency before using it for testing.21 This is not done within the simulations due to the automated procedure, but showcased in the supplementary material.
+Each combination of sample scenario and prior specification leads to a total of 168 simulation settings – 144 with a positive ADR rate and 24 controls with only background events to match.
 Each simulation scenario’s data generation and posterior estimation is repeated 100 times.
+The pWSP test does not require a prior specification, such that the number of simulation scenarios is reduced to 21 – 18 scenarios with a positive ADR rate and three controls.
+Data generation and maximum likelihood estimation are performed 800 times each.
 3.3 Specification of the HDI+ROPE test
-The HDI-ROPE test approach is specified with respect to the
+The Bayesian HDI+ROPE test approach is specified with respect to the
   • ROPE position and width,
@@
   • combination of the single test outputs of both shape parameters.
-We specify the ROPE for each shape parameter as equal-tailed interval (ETI)
+We specify the ROPE for each shape parameter $\nu$ and $\gamma$ as equal-tailed interval (ETI)
 $[q_{(1-\alpha)/2},q_{(1+\alpha)/2}]$ (9)
-based on the quantiles of the shape parameters’ prior distributions under the null-hypothesis at credibility level $\alpha$.
-We specify the posterior CI for each shape parameter either as an ETI calculated from the empirical quantiles of the shape parameters’ posterior samples at credibility level $\alpha$ or as an HDI at credibility level $\alpha$ calculated from the posterior sample.
+based on the quantiles $q$ of the shape parameters’ prior distributions under the null-hypothesis at credibility level $\alpha$.
+The posterior CI at credibility level $\alpha$ for each shape parameter is set to be an ETI (see Equation 9) calculated from the empirical quantiles of the shape parameters’ posterior samples or an HDI estimated based on the posterior sample.
 An HDI at level $\alpha$ is an interval
-[…]
-that is with boundaries such that all values inside the HDI are have higher density than values outside the HDI and simultaneously encompasses $(1-\alpha)\cdot 100\%$ of the probability mass (Kruschke,, 2015, p.87).
-To control the width of the ROPE as well as the posterior CIs, we apply credibility levels $\alpha$ ranging from $0.5,0.55,...0.95$.
-Three different combination rules of the single shape parameters’ test outcomes are defined in Table 3.
-Option 1 implements the hypothesis formulated in Equation 3.
-A signal is raised if at least one shape parameter significantly differs from one indicating a non-constant hazard function.
-Option 2 raises a signal in fewer cases, namely if at least one shape parameter’s ROPE test is positive and the other undecided.
-Option 3 raises a signal if and only if both single shape parameter tests are positive.
-The simulation study is implemented in the statistical software environment R (Team,, 2022) including the packages rstan (Stan Development Team,, 2024) for the Bayesian model setup and posterior sampling, HDInterval (Meredith and Kruschke,, 2022) for calculation of posterior CIs, and BWSPsignal (available on https://github.com/julia-dyck/BWSPsignal) containing functions specifically written for the BPgWSP test development and simulation setup.
+$HDI(\nu)=\{\nu\;|\;p_{1}(\nu)\geq w\}\\ \text{ with }w\in[0,1]\text{ such that}\int_{\nu\;|\;p_{1}(\nu)\geq w}p_{1}(\nu|t)\;d\nu=1-\alpha,$ (10)
+that is with boundaries such that all values inside the HDI have a higher density than values outside the HDI and simultaneously encompasses $(1-\alpha)\cdot 100\%$ of the probability mass.19
+To control the width of the ROPE as well as the posterior CIs, we apply credibility levels $\alpha$ ranging from 0.5,0.55,0.950.5,0.55,...0.95.
 interim result combination rule
@@
 Option 1 raises a signal if at least one parameter significantly differs from one.
-Option 2 raises a signal in fewer cases.
-Option 3 raises a signal only if both single parameter tests are positive.
+Option 2 raises a signal in fewer cases, namely when one parameter test rejects the null hypothesis $H_{0}$ and the other rejects $H_{0}$ or is undecided.
+Option 3 raises a signal only if both single parameter tests reject $H_{0}$.
+Three different combination rules of the single shape parameters’ test outcomes are defined in Table 2.
+Option 1 implements the hypothesis formulated in Equation 3.
+A signal is raised if at least one shape parameter significantly differs from one indicating a non-constant hazard function.
+Option 2 raises a signal in fewer cases, namely if the null-hypothesis is rejected for at least one shape parameter and rejected or undecided for the other.
+Option 3 raises a signal if and only if both single shape parameter tests reject the null-hypothesis.
+The pWSP test is constructed according to Sauzet and Cornelius.6 Based on the maximum likelihood estimate (MLE) the confidence interval at confidence level $\alpha$
+$[\hat{\nu}_{MLE}\pm z_{(1+\alpha)/2}\cdot sd(\hat{\nu}_{MLE})]$ (11)
+is calculated and used for significance testing for each shape parameter.
+We apply confidence levels 0.5,0.55,0.90.5,0.55,...0.9 and 0.91,0.990.91,...0.99 as well as 0.991,0.9990.991,...0.999 to take into account the relevant levels according to Sauzet et al.
+22 When shape parameters cannot be estimated (e.g.
+due to convergence issues), this is considered as no rejection of the single parameter null-hypothesis.
+A signal is raised if both parameter estimates differ significantly from the null value.
+The simulation study is implemented in the statistical software environment R 23 including the packages rstan 24 for the Bayesian model setup and posterior sampling, HDInterval 25 for calculation of posterior CIs, and BWSPsignal 26 containing functions for the BPgWSP test and simulation framework.
 3.4 Evaluation criteria for the simulation results
 The evaluation of the simulation study is performed in three steps.
-The first step consists of selecting a combination of prior distributions (see Table 2) that provides reliable posterior estimates.
-We will reject all combinations of prior distributions that lead to cases of non-convergence in the posterior estimation or that tend to higher execution times or small effective posterior sample sizes.
-According to Kruschke, (2015), a posterior sample is recommended to be used for HDI+ROPE testing if the effective sample size is at least $10000$.
-The second step includes a classification of tuning parameter combinations for the HDI+ROPE test and combination rule based on a sensitivity and specificity driven performance measure.
-The classification is restricted to prior distributions selected in step one.
-The performance of the implemented test setups (depending on ROPE and posterior CI definition, credibility level and combination rule) averaged over varying simulation scenarios (sample size of data, background rate, ADR rate, expected event-time) are measured in terms of the area under the curve (AUC) of the receiver operating characteristic graph with one threshold (Fawcett,, 2004; Lloyd,, 1998) based on equal numbers of ADR and no-ADR scenarios (and thereby equal numbers of true positives and negatives).
-The AUC is calculated using the R package ROCR (Sing et al.,, 2005).
-In the third step, the BPgWSP test with the tuning parameters chosen in step two is investigated further.
-The influence of the sample scenario on the AUC performance is presented in terms of the AUC stratified with respect to sample size, ADR rate, and expected event-time.
-Finally, the robustness to miss-specification is explored by examining the AUC when the prior belief differs from the truth because the expected event-time is incorrectly specified.
+First, prior distributions that show convergence issues, extremely high execution times, or small posterior effective sample sizes (ESS) are excluded.
+The model is considered to have not converged whenever the rstan::sampling function returns an error.
+We consider the execution time extreme if the maximum is 10 hours or more per model.
+Finally, following Kruschke’s recommendation for HDI+ROPE testing, we consider prior distributions exceeding a posterior ESS of $10,000$ or more on average as sufficient for testing.
+19, 21
+Second, the tuning parameter combinations are classified based on the test sensitivity and specificity under correct prior beliefs.
+The performance of the implemented test setups averaged over varying simulation scenarios is measured in terms of the area under the curve (AUC) of the receiver operating characteristic (ROC) graph with one threshold 27, 28 based on equal numbers of ADR and no-ADR scenarios.
+In addition to the AUC value, the ROC curve allows further investigation of the top tests by displaying the AUC depending on the false and true positive (FP, TP) rates.
+The AUC is calculated using the R package ROCR.
+29
+Third, to explore the effect of sample size, ADR rate, and expected event-time on the test performance, we present grouped AUCs of the best test.
+In case of a BPgWSP test, the robustness to miss-specification is explored by examining the AUC when the prior belief differs from the truth and the expected event-time is therefore incorrectly specified.
 3.5 Results of the simulation study
-Per prior distributional choice (see Table 2) 8400 posterior estimations (resulting from 336 simulation scenarios divided by number of distribution choices times 100 repetitions) were conducted.
-In the following, the prior distributions are compared with respect to convergence issues, execution times and effective sample sizes.
-Non-convergence in posterior estimation:
-Only few posterior estimations failed due to non-convergence of the MCMC.
-This occurred for $3.6\%$ of the posterior estimations given a fix-gam-gam prior and for $10.2\%$ of the posterior estimations when a gam-gam-gam prior was used.
-No implications for convergence issues (in form of an error message and canceled execution of posterior estimation) came up when a fix-log-log or log-log-log prior was specified.
-As they seem to provide smooth estimation without the need to adjust the MCMC sampler, we suggest to use a fix-log-log or a log-log-log prior distribution.
-Execution times:
-Figure 5: Boxplots of running times split with respect to prior distributional choice.
-The graph is cut on the y-axis at 210 minutes (3h) although the boxplot for the gam-gam-gam prior reaches as maximum at about 45 hours.
+We compare the prior distributions with respect to convergence issues, execution times, and ESSs. There were no convergence issues.
+Figure 5: Boxplot comparing execution times in minutes for two prior distributions, gamma (pgw-$g$) and lognormal (pgw-$l$), with lognormal yielding higher maximum run times of up to 3 h, but both prior distributions producing similar median times of 16 Minutes with gamma and 17 Minutes with lognormal prior.
+Figure 6: Boxplots of estimated effective sample sizes (n_eff) split with respect to prior distributional choices gamma (pgw-$g$) and lognormal (pgw-$l$) for parameter $\nu$ (shape1 in red) and $\gamma$ (shape2 in blue).
+The recommended threshold for the effective sample size is marked with a horizontal dashed line at 10 00010\,000.
 A boxplot of the execution times for posterior estimation stratified with respect to the prior distribution is shown in Figure 6.
-All distribution choices lead to acceptable median execution times of 8 min for fix-log-log models, 13 min for log-log-log models, 10 min for fix-gam-gam models, and 40 min for gam-gam-gam models.
-While the maximal running times that occurred during the simulation stayed under about 1 h given a fix-log-log, 2 h given a fix-gam-gam, or 3 h given a log-log-log distribution, the gam-gam-gam prior produced execution times of up to 45 hours.
-Effective sample sizes:
-Figure 6: Boxplots of estimated effective sample size split with respect to prior distributional choice for parameter $\nu$ and $\gamma$.
-In $73\%$ of the posterior estimations for $\nu$ and $55\%$ of the posterior estimations for $\gamma$ does the log-log-log model generate a posterior sample with an effective sample size $>10000$.
-The fix-log-log, fix-gam-gam and gam-gam-gam prior distributions provide only $\approx 25-28\%$ simulations for $\nu$ and $21-26\%$ for $\gamma$ fulfilling the rule of thumb (see Figure 6).
-This makes the log-log-log prior preferable compared to the alternative prior distributions.
-The log-log-log distribution provides the best results in terms of convergence of the posterior estimation, adequate running times and sufficiently high effective posterior sample size in most cases.
-Consequently, we now proceed to evaluate the simulation results of the the log-log-log prior distribution.
-Ranking of test specifications based on the AUC:
-We rank all combinations of ROPE and posterior CI specifications, credibility levels and combination rules based on the AUC averaged over all sample scenarios given a correct prior belief.
-Table 4 shows the top ten test specifications in terms of AUC.
-The highest AUC value of $0.89$ is achieved with an $80\%$ Lognormal ETI as ROPE, an $80\%$ posterior HDI as posterior CI and option 2 for the combination rule (see Table 3).
-rank ROPE posterior CI combination rule AUC
-1 80% ETI 80% HDI option 2 0.890
-2 80% ETI 80% ETI option 2 0.873
-3 80% ETI 80% ETI option 3 0.838
-4 80% ETI 80% HDI option 3 0.834
-5 75% ETI 75% HDI option 3 0.823
-6 75% ETI 75% ETI option 3 0.802
-7 85% ETI 85% HDI option 2 0.693
-8 85% ETI 85% ETI option 2 0.691
-9 90% ETI 90% HDI option 1 0.653
-10 85% ETI 85% HDI option 1 0.643
-Table 4: Ranking of top ten BPgWSP test specifications according to AUC averaged over all sample scenarios given a correct prior belief.
-Effect of simulation scenario on AUC performance:
-The influence of the sample size, the proportion of ADR events and the expected event-time of the ADR (with correctly specified prior belief) is summarized in Table 5.
-The BPgWSP test works better with increasing sample size reaching an AUC of $98.9\%$ given a sample size of 5000.
-An increase in the proportion of ADR events seems to improve detectability of the ADR very slightly.
-The test performance is higher for ADRs occurring around the first or third quarter of the observation period with an AUC of $94\%$, but still provides an AUC of $78\%$ for ADRs occurring in the middle of the observation period.
+The models produce median execution times of 16 Minutes with gamma priors, and 17 Minutes with lognormal priors.
+The maximal running times stayed under 1 h 40 min given a gamma prior and under 3h given a lognormal prior.
+Gamma priors lead to ESSs of >10 000>10\,000 in $85\%$ of the posterior estimations for $\nu$, and $68\%$ for $\gamma$.
+For lognormal priors these number are $73\%$ for $\nu$ and $55\%$ for $\gamma$ (see Figure 6).
+Based on execution time and ESS results, both prior distribution choices are acceptable for the modeling part of the BPgWSP.
+Figure 7: Roc curves representing the top one (left) and top ten (right) BPgWSP tests’ performances.
+rank prior ROPE posterior CI combination AUC FP rate TP rate
+distribution rule
+1 lognormal 80% ETI 80% HDI option 2 0.900 0.173 0.974
+2 lognormal 80% ETI 80% ETI option 2 0.876 0.226 0.977
+3 lognormal 80% ETI 80% ETI option 3 0.844 0.006 0.694
+4 lognormal 80% ETI 80% HDI option 3 0.837 0.001 0.675
+5 lognormal 75% ETI 75% HDI option 3 0.823 0.326 0.971
+6 gamma 93% ETI 93% HDI option 2 0.810 0.356 0.976
+7 lognormal 75% ETI 75% ETI option 3 0.808 0.364 0.980
+8 gamma 93% ETI 93% ETI option 3 0.804 0.003 0.611
+9 gamma 93% ETI 93% ETI option 2 0.803 0.371 0.978
+10 gamma 93% ETI 93% HDI option 3 0.797 0.002 0.597
+Table 3: Top ten BPgWSP and pWSP test specifications ranked based on AUC. AUC, FP rate and TP rate are averaged over all sample scenarios given a correct prior belief.
+Table 3 shows the top ten test specifications in terms of AUC along with the corresponding FP and TP rates.
+Figure 7 displays the ROC curves of the top one (left) and top ten (right) test specifications.
+The highest AUC value of $0.9$ is achieved with an $80\%$ lognormal ETI as ROPE, an $80\%$ posterior HDI as posterior CI and option 2 for the combination rule.
+The corresponding ROC curve exhibits a TP rate of $0.97$ and a FP rate of $0.17$, indicating a trade-off that prioritizes the detection of TPs.
+The best pWSP test (frequentist) with confidence level of $55\%$ is ranked in 11th place with an AUC of $0.75$ under the given simulation conditions.
+The influence of the sample size $N$, the proportion of ADR events and the expected event-time of the ADR on the AUC under a correctly specified prior belief is summarized in Table 4.
+The BPgWSP test works better with increasing $N$ reaching an AUC of $0.978$ given an $N$ of 5000.
+An increase in the proportion of ADR events from $5\%$ to $10\%$ of $N$ increases the AUC very slightly from $0.888$ to $0.913$.
+The test performance is higher for ADRs occurring around the first or third quarter of the OP with an AUC of $0.969$ or $0.946$, respectively, but still provides an AUC of $0.786$ for ADRs occurring in the middle of the OP.
 grouping factors AUC
-sample size $N$
-$500$ 0.749
-$3000$ 0.932
-$5000$ 0.989
+sample size NN
+$500$ 0.789
+$3000$ 0.933
+$5000$ 0.978
 proportion of ADR events
-$5\%$ of $N$ 0.881
-$10\%$ of $N$ 0.899
+$5\%$ of NN 0.888
+$10\%$ of NN 0.913
 expected time of ADR events
-1st quarter of observation period 0.946
-2nd quarter of observation period 0.782
-3rd quarter of observation period 0.942
-Table 5: Average AUC performance of the BPgWSP test (prior distribution: log-log-log, ROPE: $80\%$ Lognormal confidence interval, credibility interval: $80\%$ highest posterior density region, combination rule option 2) grouped by sample size, proportion of ADR events, and expected time of ADR events.
-Robustness to miss-specification:
-If the specification of the expected event-time is off by one quarter of the observation time, the AUC reduces to $78\%$ on average.
-If the a priori expected event-time is off by two quarters, the AUC performance drops to $47\%$.
-An AUC performance of $79\%$ is provided in case no ADR is suspected a priori.
-For application of the BPgWSP test to a range of drug-AE pairs (in an automated manner), we recommend to set up the tuning parameters as follows: an $80\%$ Lognormal ETI for the ROPE, an $80\%$ posterior HDI for the posterior CI and option 2 for the combination rule.
-In the next section, the developed method will be applied to real data for demonstration purposes using the recommended tuning for the BPgWSP test.
+1st quarter of observation period 0.969
+2nd quarter of observation period 0.786
+3rd quarter of observation period 0.946
+Table 4: Average AUC performance of the BPgWSP test with lognormal prior distribution, $80\%$ lognormal confidence interval as ROPE, $80\%$ HDI as posterior CI, and combination rule option 2 grouped by sample size, proportion of ADR events, and expected time of ADR events given a correct prior belief.
+If the specification of the expected event-time is off by one quarter of the observation time, the AUC reduces to $0.783$ on average.
+If the a priori expected event-time is off by two quarters, the AUC performance drops to $0.489$.
+An AUC performance of $0.787$ is provided in case no ADR is suspected a priori (see Table 5).
 type of prior belief AUC
-correct specification 0.890
-one quarter off 0.780
-two quarters off 0.469
-no ADR assumed 0.788
+correct specification 0.900
+one quarter off 0.783
+two quarters off 0.489
+no ADR assumed 0.787
 Table 5: AUC performance given a correct or incorrect prior belief of different types about the expected event-time.
+For application, we recommend setting the tuning parameters as follows: an $80\%$ lognormal ETI for the ROPE, an $80\%$ posterior HDI for the posterior CI and option 2 for the combination rule.
 4 Case study
-The purpose of this case study is to illustrate the BPgWSP test derived in this work with a real dataset.
-Conclusions from this case study should not be used to inform clinical practice.
-The dataset was originally created from the Health Improvement Network (THIN) database (THIN,, 2024) to assess the association between the prescription of bisphosphonates and the occurrence of Carpal tunnel syndrome in menopausal Women (Carvajal et al.,, 2016).
-The THIN database contains primary care data from patients in the UK.
+The purpose of this case study is to illustrate the BPgWSP test with a real dataset.
+Conclusions from this section should not be used to inform clinical practice.
+The dataset was originally created from the Health Improvement Network (THIN) database 30 to assess the association between the prescription of bisphosphonates and the occurrence of carpal tunnel syndrome in menopausal women.31 The THIN database contains primary care data from patients in the UK.
 The BPgWSP test is applied to investigate the association between bisphosphonates and four AEs occurring in the population.
 In addition to the exposed cohort, the dataset contains patients not exposed to bisphosphonates (control group).
-A control group is not required to apply the BPgWSP test.
-Nevertheless, it provides us with the opportunity to explore the type one error, i.e.
-the test results for the control group.
-The data preparation has been described in Sauzet et al., (2013).
-To facilitate the description of the data, we recall the descriptive statistics provided in Sauzet et al., (2013).
-All steps are conducted for the exposed and the control groups.
-We present the sample size and number of events as well as the mean event-time (and standard deviation).
-A histogram of event-times gives insights into the distribution over time.
-A smoothed non-parametric estimate of the hazard function provides information on the trend of risk for an AE to occur over time.
-The BPgWSP test is applied with the recommended tuning parameter resulting from the simulation study (log-log-log priors, 80% Lognormal confidence interval for the ROPE, 80% HDI for the posterior CI, option 2 for the combination rule).
-Selectively, we compare our test outcomes to the test results obtained with the non-Bayesian approaches illustrated in Sauzet et al., (2013) and Sauzet and Cornelius, (2022).
+A control group is not required to apply the BPgWSP test, but allows to explore potential type I errors.
+The data is described in more detail in Sauzet et al.
+32
+We illustrate the BPgWSP test using four different prior beliefs for the lognormal parameters reflecting an a priori expected ADR event-time at the first, second, or third quartile of the OP, or none.
+Given the lack of prior knowledge, we choose priors based on visualizing the corresponding hazard.
+Otherwise, we used the same model set-up as in the simulation study.
+If the ESS is lower than 10 00010\,000 or the scale reduction factor21 $\hat{R}$ far from one, a larger posterior sample is generated.
+The BPgWSP test is applied with the recommended tuning of the simulation study.
+The R code and results are stored on https://osf.io/p8hvr/files/osfstorage. See the supplementary material for a more detailed description of the analysis.
 4.1 Headache
-Figure 7: Graphics for exploratory analysis of the occurrence of headache over 21 days. The graphic derived is based on Sauzet et al., (2013, Fig. 1).
+Figure 8: Graphics for exploratory analysis of the occurrence of headache over 21 days. The graphic derived is based on Sauzet et al., Fig. 1. 32
 Exploratory analysis:
@@
 In this time, 21 among 19791 patients $(0.11\%)$ exposed to bisphosponates experienced headaches.
-The day of occurrence is on average $14\;(5.9)$.
-The histogram of event-times (Figure 7 a.1) shows that the number of events is slightly higher in the second half of the observation period.
+The day of occurrence is on average 14(sd=5.9)14\;(sd=5.9).
+The histogram of event-times (Figure 8 a.1) shows that the number of events is slightly higher in the second half of the OP.
 The non-parametric smooth hazard estimate (Figure 8 b.1) indicates a more pronounced increase of risk in the last seven days.
@@
 The hazard estimate is almost horizontal (see Figure 8 b.2).
-prior belief about prior means test outcome test outcome
+prior belief about a priori $E[t]$ prior means test outcome test outcome
 expected event-time $(\bar{\theta}_{0},\bar{\nu}_{0},\bar{\gamma}_{0})$ (Exposed) (Control)
-none $(1,1,1)$ signal -
-1st quarter of OP (5.25 days) $(2.5,0.5,1)$ signal -
-2nd quarter of OP (10.5 days) $(3,3,5)$ signal -
-3rd quarter of OP (15.75 days) $(18,5,1)$ signal -
-Table 7: Test outcomes under various prior beliefs for headache based on exposed and control group.
-Figure 8: Prior distribution and posterior sample histogram obtained from Bayesian model under the prior belief that headache is no ADR.
-The test results in row one of Table 7 are concluded from the alignment of the ROPE and HDIs and combination according to combination rule 2.
+none - $(1,1,1)$ - -
+1st quarter of OP 5.25 days $(2.5,0.5,1)$ - -
+2nd quarter of OP 10.5 days $(3,3,5)$ signal -
+3rd quarter of OP 15.75 days $(18,5,1)$ signal -
+Table 6: Test outcomes under various prior beliefs for headache based on exposed and control group. Prior standard deviation is set to 10 for all parameters in all prior belief cases.
 Signal detection test:
-The prior specifications and results of the BPgWSP test are shown in Table 7.
-The signal detection tool is applied to the data censored at 21 days under the prior beliefs that the symptom occurrence is expected at the first, second or third quarter of the observation period, as well as a prior belief of no association.
-Independent of the prior specification including the assumption of no association, a signal is consistently raised for the exposed group and no signal for the control group.
-This is in line with the results of the non-Bayesian approach.
-The intermediate shape parameters’ HDI+ROPE test outcomes are illustrated in Figure 8 for the exposed and control group under the prior belief of no association.
-In the case of the exposed cohort, both single parameter tests reject the null-hypothesis as the HDI does not lie within the ROPE) and combination scheme option 2 concludes that a signal is raised.
-In contrast, the control group’s HDI+ROPE test for $\gamma$ returns a rejection of the null-hypothesis while the HDI+ROPE test for $\nu$ returns an acceptance.
-According to the combination scheme option 2, no signal is raised.
+A signal is raised for the exposed group under the prior belief that ADR occurs towards the middle or end of the OP.
+No signal is raised under the prior of no ADR or ADR at the beginning of the OP (see Table 6).
+For the control group, no signal is raised independent of the prior specification.
+Headache is a known ADR of bisphosphonates.
+33 The BPgWSP test detects a signal under the prior belief of ADR occurrence towards the middle or end of the OP.
+These beliefs presumably amplify the signal in the data already indicating a higher hazard rate towards the end of the OP (see Figure 8 b.1) while the other prior settings might fail to detect a signal due to the contradictory prior and data information.
 4.2 Musculoskeletal pain
-Figure 9: Graphics for exploratory analysis of the occurrence of musculoskeletal pain over one year. The graphic derived is based on Sauzet et al., (2013, Fig. 2).
+Figure 9: Graphics for exploratory analysis of the occurrence of musculoskeletal pain over one year. The graphic derived is based on Sauzet et al., Fig. 2. 32
 Exploratory analysis:
-The data is censored after one year in which 207 of 19777 ($1\%$) exposed patients reported musculoskeletal pain.
-The average event-time to occurrence is 160 days $(99)$.
-The histogram of event-times (Figure 9 a.1) shows a slightly higher density in the first half of the observation period.
+The data is censored after one year in which 344 of 19777 ($1.7\%$) exposed patients reported musculoskeletal pain.
+The average event-time is 160 days $(sd=99)$.
+The histogram of event-times (Figure 9 a.1) shows a slightly higher density in the first half of the OP.
 The smoothed hazard estimate (Figure 9 b.1) indicates a slight increase in risk during the first four months followed by a decrease.
-In the control group musculoskeletal pain was reported 207 times from 39634 patients ($0.5\%$).
-According to the histogram (Figure 9 a.2) slightly more cases are reported in the first and third month but evenly distributed for the rest of the time.
+In the control group musculoskeletal pain was reported 366 times from 39634 patients ($0.9\%$).
+According to the histogram (Figure 9 a.2) slightly more cases are reported in the first and third month and evenly distributed for the rest of the time.
 The hazard estimate (Figure 9 b.2) looks almost constant.
-prior belief about prior means test outcome test outcome
+prior belief about a priori $E[t]$ prior means test outcome test outcome
 expected event-time $(\bar{\theta}_{0},\bar{\nu}_{0},\bar{\gamma}_{0})$ (Exposed) (Control)
-none $(1,1,1)$ signal -
-1st quarter of OP (91 days) $(1,0.207,1)$ signal -
-2nd quarter of OP (183 days) $(20,5.5,14)$ signal -
-3rd quarter of OP (274 days) $(300,4,1)$
-Table 8: Test outcomes under various prior beliefs for musculoskeletal pain based on exposed and control group.
+none - $(1,1,1)$ signal -
+1st quarter of OP 91 days $(1,0.207,1)$ signal -
+2nd quarter of OP 183 days $(20,5.5,14)$ signal -
+3rd quarter of OP 274 days $(300,4,1)$ - -
+Table 7: Test outcomes under various prior beliefs for musculoskeletal pain based on exposed and control group. Prior standard deviation is set to 10 for all parameters in all prior belief cases.
 Signal detection test:
-The results of the BPgWSP test applied to the data censored at 365 days under the prior beliefs of symptom occurrence at the first, second or third quarter of the observation period or of no association are shown in Table 8.
 A signal is raised for the exposed group under all priors except a prior belief around the third quarter (day 274).
-For the control group, no signal is raised regardless of the prior choice.
-In the case of musculoskeletal pain, the test results depend on the prior specification.
-Here we have a long observation period and a small number of cases which potentially decreases the performance of the BPgWSP test.
-Given that most cases occur at the beginning of the observation period the prior specification for the last observation quarter does not lead to a signal.
-This presumably occurs because data and prior belief provide contrasting information.
+For the control group, no signal is raised regardless of the prior choice (see Table 7).
+Musculoskeletal pain is a recognized ADR that can occur anytime after the beginning of bisphosphonate intake.34 Here, the test result depends on the prior specification.
+A long OP and a small number of cases potentially decrease the performance of the BPgWSP test when the prior contrasts sharply with the data, as is the case under the end-prior here.
+Moreover, the result of no signal for the exposed group that was based on the model that was not rerun might indicate the need for more iterations, even when $\hat{R}$ and ESS suggest convergence.
 4.3 Alopecia
-Figure 10: Graphics for exploratory analysis of the occurrence of alopecia over three years. The graphic derived is based on Sauzet et al., (2013, Fig. 3).
+Figure 10: Graphics for exploratory analysis of the occurrence of alopecia over three years. The graphic derived is based on Sauzet et al., Fig. 3. 32
 Exploratory analysis:
-Over three years there have been 144 reports of alopecia in the group of 19812 patients (0.73%) exposed to bisphosphonates.
-The mean event-time is on day 424 (285).
-According to the histogram (Figure 10 1.a) events tend to take place slightly more often between the fifth and 13th month which roughly corresponds to the period around the first or second quarter of the observation period.
-Regarding the estimated hazard function the risk rises and falls within the first two years and stabilizes after that.
+Over three years, there have been 144 reports of alopecia in the group of 19812 patients (0.73%) exposed to bisphosphonates.
+The mean event-time is day 424 $(sd=285)$.
+According to the histogram (Figure 10 1.a), events take place slightly more often between the fifth and 13th month which roughly corresponds to the period around the first or second quarter of the OP.
+Regarding the estimated hazard function, the risk rises and falls within the first two years and stabilizes after that.
 There are 180 alopecia cases among 39634 control patients ($0.45\%$).
-The event-times are slightly more dense in the first half of the observation window.
+The event-times are slightly denser in the first half of the observation window.
 The estimated hazard (Figure 10 2.b) has an almost flat trend.
+prior belief about a priori $E[t]$ prior means test outcome test outcome
+expected event-time $(\bar{\theta}_{0},\bar{\nu}_{0},\bar{\gamma}_{0})$ (Exposed) (Control)
+none - $(1,1,1)$ signal -
+1st quarter of OP 274 days $(200,0.63,1)$ signal -
+2nd quarter of OP 548 days $(30,3.2,10)$ signal -
+3rd quarter of OP 821 days $(700,14,4)$ - -
+Table 8: Test results under various prior beliefs for alopecia based on exposed and control group. Prior standard deviation is set to 10 for all parameters in all prior belief cases.
 Signal detection test:
-prior belief about prior means test outcome test outcome
-expected event-time $(\bar{\theta}_{0},\bar{\nu}_{0},\bar{\gamma}_{0})$ (Exposed) (Control)
-none $(1,1,1)$ signal -
-1st quarter of OP (274 days) $(200,0.63,1)$ signal -
-2nd quarter of OP (548 days) $(30,3.2,10)$ signal -
-3rd quarter of OP (821 days) $(700,14,4)$
-Table 9: Test results under various prior beliefs for alopecia based on exposed and control group.
-Table 9 shows the BPgWSP test results for alopecia.
-Under the prior assumptions of no association, ADR occurrence in the beginning or in the middle of the observation period the test identifies a temporal pattern for the exposed group and none for the control group.
-For both groups, no signal is raised under the prior belief that alopecia risk increases at the end of the observation period.
-Alopecia is a rare event.
-In the non-Bayesian approach in Sauzet et al., (2013) an ambiguous signal was given if the data was censored after one year only whereas here, we get a consistent signal if the prior belief does not expect cases to occur toward the three-year mark.
-For the control group, no signal is raised indicating that the signal raised for cases is probably not due to noise.
+Under the prior assumptions of no association, ADR occurrence in the beginning or in the middle of the OP the test identifies a temporal pattern for the exposed group and none for the control group.
+For both groups, no signal is raised under the prior belief that alopecia risk increases at the end of the OP (see Table 8).
+Alopecia is a rare event not certainly identified as an ADR of bisphosphonates at the current state of research to our knowledge.
+The same uncertainty is reflected in the BPgWSP test results.
+Under an ADR-suspecting prior, no signal is raised for the control group and a signal flagged for the exposed group indicating that the signal might not be raised due to noise.
+Even with larger posterior samples under the none-prior, no signal was identified.
+This indicates that more iterations do not necessarily lead to a higher signal rate.
 4.4 Carpal tunnel syndrome
-Figure 11: Graphics for exploratory analysis of the occurrence of carpal tunnel syndrome over three years. The graphic derived is based on Sauzet et al., (2013, Fig. 4).
+Figure 11: Graphics for exploratory analysis of the occurrence of carpal tunnel syndrome over three years. The graphic derived is based on Sauzet et al., Fig. 4. 32
 Exploratory Analysis:
-Within the observation period of three years, 192 among 19817 exposed patients ($0.97\%$) experienced carpal tunnel syndrome.
-The day of occurrence is on average $429\;(5.9)$ which is approximately one year after drug prescription.
-The histogram of event-times (Figure 11 a.1) reveals that events very slightly decrease over the scope of the observation period.
+Within the OP of three years, 192 among 19817 exposed patients ($0.97\%$) experienced carpal tunnel syndrome.
+The day of occurrence is on average 429(sd=290)429\;(sd=290) which is approximately one year after drug prescription.
+The histogram of event-times (Figure 11 a.1) reveals that events very slightly decrease over the scope of the OP.
 The hazard estimate (Figure 11 b.1) indicates a slightly higher constant risk in the first 18 months and a decrease after that.
-Within the control group, carpal tunnel syndrome occurred in 244 among 39389 patients ($0.62\%$).
+Within the control group, carpal tunnel syndrome occurred for 244 among 39389 patients ($0.62\%$).
 The distribution of events looks similar to the exposed group’s plot (see Figure 8 a.2).
@@
 Signal detection test:
-prior belief about prior means test outcome test outcome
+In two out of four prior specifications, namely when the prior belief is that the event is to be expected around the first quarter or middle of the OP, a signal is raised for both the control and exposed group.
+In the two other cases, a signal is raised for neither group (see Table 9).
+Carpal tunnel syndrome is not a recognized ADR to our knowledge.
+The ambiguous test results assigning signals to exposed and control group under a subset of prior specifications are probably returned due to the confounding effect of the age of the participants for which the carpal tunnel syndrome is a common condition.
+This suggests that sample stratification might be an important preparation before BPgWSP testing in some applications.
+prior belief about a priori $E[t]$ prior means test outcome test outcome
 expected event-time $(\bar{\theta}_{0},\bar{\nu}_{0},\bar{\gamma}_{0})$ (Exposed) (Control)
-none $(1,1,1)$
-1st quarter of OP (274 days) $(200,0.63,1)$ signal signal
-2nd quarter of OP (548 days) $(30,3.2,10)$ signal signal
-3rd quarter of OP (821 days) $(700,14,4)$
-Table 10: Test outcomes under various prior beliefs for carpal tunnel syndrome based on exposed and control group.
-The BPgWSP test results for carpal tunnel syndrome are presented in Table 10.
-In two out of four prior specifications, namely when the prior belief is that the event is to be expected around the first quarter or middle of the observation period, a signal is raised for both control and exposed group.
-In the two other cases, a signal is raised for neither group.
-The ambiguous test results occurring here and in Sauzet et al., (2013) where no signal was raised for the exposed group and a signal was raised for the control group when censored after 637 days are probably returned due to the confounding effect of the age of the participants for which the carpal tunnel syndrome is a common condition.
+none - $(1,1,1)$ - -
+1st quarter of OP 274 days $(200,0.63,1)$ signal signal
+2nd quarter of OP 548 days $(30,3.2,10)$ signal signal
+3rd quarter of OP 821 days $(700,14,4)$ - -
+Table 9: Test outcomes under various prior beliefs for carpal tunnel syndrome based on exposed and control group. Prior standard deviation is set to 10 for all parameters in all prior belief cases.
 5 Discussion
@@
 The defined test was tuned based on a simulation study which led to a one-fits-all approach.
-Consequently, we applied the test with the recommended tuning to real data.
-The underlying assumption of the BPgWSP test is that the hazard should be constant if there is no drug-AE association.
-This assumption also underlies the work of Cornelius et al., (2012) and the extension by Sauzet and Cornelius, (2022) in the frequentist framework.
-While this heuristic assumption can be refuted due to various confounding factors, we argue that it is a simplified model is good enough for flagging possible associations because finding causal association is not the aim of the procedure.
+An important aspect of our proposed BPgWSP test is the role of prior information in shaping its performance.
+The BPgWSP test performs very well under correct prior assumptions and outperforms the frequentist approach.
+If the prior specification is wrong, the performance of the test decreases considerably which is also reflected in the case study.
+We recommend using the BPgWSP test only in case of certainty about prior information, and otherwise, use the frequentist approach.
+In light of these results, correct prior elicitation is of importance.
+It is advised to follow procedures recommended by others.
+See O’Hagan for an introduction.35
 The simulation study enabled an initial assessment of the test performance and the conclusion of a recommended BPgWSP test tuning for application in an automated one-fits-all manner to a multitude of drug-AE combinations.
-Arguably, the amount of combinations of simulation scenarios and prior specifications considered within the frame of the simulation study is too limited to draw general conclusions on tuning parameters and only give advise for tuning choices.
-For instance only one censoring time (365) was considered for data generation and only one standard deviation (10) was considered for the prior belief specification such that insights on the expected performance in case of alternative specifications cannot be drawn.
-This was due to the practicality of presenting the methods.
-In general however, we recommend to tune the BPgWSP test with an additional simulation study adjusted for the circumstances of interest.
-To do so, we provide the R package BWSPsignal available on https://github.com/julia-dyck/BWSPsignal such that new values to specify the data generating characteristics (see sec.
-3.1), model specifications (see sec.
-3.2) and tuning parameters for the HDI+ROPE test setup (see sec.
-3.3) can be easily evaluated.
-In the case study, we applied the BPgWSP test to real data on bisphosphonates with adverse events known to be ADRs caused by the drug (headache, musculoskeletal pain) and adverse events not identified as ADR (alopecia, carpal tunnel syndrome).
-The BPgWSP test tuned based on the simulation results was able to raise a signal under varying prior beliefs for the known ADRs.
-It returned ambiguous test results for adverse events not categorized as ADRs under some prior belief specifications.
-The performance of the BPgWSP test depends on whether the observation period encompasses the expected event-time and on the rareness of the adverse event.
-This is an additional argument for the need of more differentiated and context specific tuning of the BPgWSP test.
-Simulations representing the data scenarios better could improve the test performance in the case study applications.
-The data used for the case study is a historical dataset derived from the THIN database (THIN,, 2024).
-It would be desirable to test the newly developed methods with a larger set of drug-AE pairs.
-However, EHR data are not easily available for method development due todata privacy regulations and costs.
-There are attempts to create synthetic, yet realistic datasets.
-An example is the SYNTHEA database (Walonoski et al.,, 2017).
-Even though the SYNTHEA database is based on a complex data-generating process including modules for a set of health conditions and medications that occur with probabilities and prevalences reflecting reality, the generating process does not include AEs.
-We think that further development of SYNTHEA and similarly structured databases needs to be supported even more extensively to facilitate the development of statistical methods for pharmacoepidemiology and other fields.
-6 Conclusion
-Signal detection is an important component of pharmacovigilance to monitor drugs post market release.
-To improve further development of statistical signal detection methods, we combine two characteristics of existing approaches in the literature, the use of longitudinal data from EHRs and the incorporation of prior knowledge about the drug under investigation, in one test.
-The Bayesian WSP test is based on the constant-hazard assumption and consists of fitting a Bayesian PgW model with a prior distribution for the PgW parameters reflecting the prior belief about the drug-AE relationship, as well as a tuned HDI+ROPE test based on the posterior shape parameters.
-We showed how to tune the parameters specif using a simulation study, and illustrated it on data about bisphosphonate prescriptions.
+In general, however, we recommend tuning the BPgWSP test with an additional simulation study.
+For this purpose, we developed the R package BWSPsignal.26
 Data availability statement
+Simulation study output as well as the R package BWSPsignal and R scripts for reproducibility are available on https://github.com/julia-dyck/BWSPsignal.
+The data used in the case study section was originally provided from the THIN 30 database. Please contact the corresponding author upon reasonable request.
 Funding